### v3

# Ideal transition systems

Abstract
We study an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring. This method starts from a given set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish a particular transition recurrence among the chosen ideals. We describe explicit constructions of such systems in two motivating cases—namely, for the ideals of matrix Schubert varieties and their skew-symmetric analogues. Despite many formal similarities with these examples, for the symmetric versions of matrix Schubert varieties, it is an open problem to construct the same kind of transition system. We present several conjectures that would follow from such a construction, while also discussing the special obstructions arising in the symmetric case.
1 Introduction
Fix a positive integer $n$ and let $\mathsf{Mat}_{n\times n}$ be the variety of $n\times n$ matrices over an algebraically closed field $\mathbb{K}$. Write $\mathsf{B}_{n}\subset\mathsf{Mat}_{n\times n}$ for the group of invertible lower triangular matrices in this set.
Recall that a partial permutation matrix is a binary matrix with at most one nonzero entry in each row and column. A version of Gaussian elimination shows that every $M\in\mathsf{Mat}_{n\times n}$ can be written as $b_{1}wb_{2}^{\top}$ for $b_{1},b_{2}\in\mathsf{B}_{n}$ and a unique partial permutation matrix $w$. Equivalently, the $n\times n$ partial permutation matrices are representatives for the distinct $\mathsf{B}_{n}\times\mathsf{B}_{n}$-orbits on $\mathsf{Mat}_{n\times n}$. These orbits are called matrix Schubert cells $\mathsf{MSC}_{w}$, and their Zariski closures are called matrix Schubert varieties $\mathsf{MSV}_{w}$. We can express these sets using rank conditions as
$\mathsf{MSC}_{w}$ $=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\},$ (1.1)
$\mathsf{MSV}_{w}$ $=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\}.$
Here we write $n]=\{1,2,\ldots,n\}$, and if $R,C\subseteq[n]$ then $M_{RC}$ denotes the submatrix of $M$ in rows $R$ and columns $C$. Thus, $M_{[i][j]}$ is the upper-left $i\times j$ corner of $M$.
Let $\mathcal{X}=[x_{ij}]_{i,j\in[n]}$ be a generic matrix of indeterminates and let $\mathbb{K}[\mathsf{Mat}_{n\times n}]=\mathbb{K}[x_{ij}:i,j\in[n]]$ be the coordinate ring of $\mathsf{Mat}_{n\times n}$. For a partial permutation matrix $w$, let $I_{w}$ be the ideal in $\mathbb{K}[\mathsf{Mat}_{n\times n}]$ generated by all minors of size $\operatorname{rank}w_{[i][j]}+1$ in $\mathcal{X}_{[i][j]}$ for $i,j\in[n]$. Since $\operatorname{rank}M\leq r$ if and only if all $(r+1)\times(r+1)$ minors of $M$ vanish, the affine variety $\mathsf{MSV}_{w}$ is exactly the zero set of $I_{w}$.
Fulton [13] showed that the equations setting these minors to zero also define $\mathsf{MSV}_{w}$ scheme-theoretically, i.e., that $I_{w}$ is a radical ideal (in fact, it is prime). Knutson and Miller [25] proved that this generating set of minors is actually a Gröbner basis for $I_{w}$ under an appropriate term order. This fact allows for convenient computation of certain cohomological invariants associated to $I_{w}$, and can also be used to derive Fulton’s result.
In [31], we reproved these results using a new inductive approach. The key step in this approach was establishing a transition recurrence of the form
$I_{w}+\langle x_{ij}\rangle=\bigcap_{v\in\Phi}I_{v}$ (1.2)
and using it to deduce properties of $I_{w}$ from inductively known properties of the ideals $I_{v}$. Here $x_{ij}$ is a particular variable not contained in $I_{w}$, $\langle x_{ij}\rangle$ is the principal ideal that $x_{ij}$ generates, and $\Phi$ is a certain set of partial permutations.
There is a geometric reason for such recurrences to exist. If $w$ and $y$ are two partial permutation matrices, then $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is a closed $\mathsf{B}_{n}$-stable set and therefore decomposes as a union $\bigcup_{v\in\Phi}\mathsf{MSV}_{v}$. A result of Ramanathan [34] implies that the scheme $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is actually a reduced union of matrix Schubert varieties, so the stronger equality $I_{w}+I_{y}=\bigcap_{v\in\Phi}I_{v}$ also holds. The transition recurrence then follows by an appropriate choice of $y$.
Example 1.1.
Let $w=w_{1}w_{2}w_{3}=132=\left[\begin{smallmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{smallmatrix}\right]$, where we identify $w$ with the permutation matrix having $1$’s in positions $(i,w_{i})$. Then every rank condition $\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ is vacuous except when $i=j=1$, giving
$\mathsf{MSV}_{132}=\left\{A\in\mathsf{Mat}_{3\times 3}:\operatorname{rank}A_{[2][2]}\leq 1\right\}\qquad\text{and}\qquad I_{132}=\langle u_{21}u_{12}-u_{11}u_{22}\rangle.$
The transition recurrence in this case reads as
$I_{132}+\langle u_{11}\rangle=\langle u_{11},u_{21}u_{12}\rangle=\langle u_{11},u_{21}\rangle\cap\langle u_{11},u_{12}\rangle=I_{231}\cap I_{312}.$
The situation is much the same if we replace $\mathsf{Mat}_{n\times n}$ with the subset $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ of skew-symmetric $n\times n$ matrices over $\mathbb{K}$, now considering the orbits of the $\mathsf{B}_{n}$-action $b\cdot M=bMb^{\top}$. The corresponding $\mathsf{B}_{n}$-orbits $\mathsf{MSC}^{\mathrm{ss}}_{w}$ in $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ and their closures $\mathsf{MSV}^{\mathrm{ss}}_{w}$ are defined by the same rank conditions as in (1.1), but now the orbit representatives $w$ are skew-symmetric $\{0,1,-1\}$-matrices with at most one nonzero entry in each row and column.
In our previous work [31], we constructed Gröbner bases for the prime ideals of the skew-symmetric matrix Schubert varieties $\mathsf{MSV}^{\mathrm{ss}}_{w}$. These generating sets consist of Pfaffians of principal submatrices of a generic skew-symmetric matrix, and are more complicated than the minors generating the ideals $I_{w}$ above. However, the proof technique based on transition recurrences is the same as in the classical case. It would be interesting to know if there is some geometric result similar to [34] that explains these skew-symmetric transition recurrences; at present, this is an open question.
The first, partially expository goal of this article is to present a streamlined version of the proof techniques just described. Specifically, we formalize an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring, in terms of what we call transition systems. Such a system is composed of a set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish transition recurrences analogous to (1.2) among the chosen ideals. Section 3 develops the general setup and nontrivial properties of transition systems, following some preliminaries in Section 2. Our motivating examples appear in Sections 4 and 5, where we present transition systems that compute Gröbner bases for the matrix Schubert varieties $\mathsf{MSV}_{w}$ and $\mathsf{MSV}^{\mathrm{ss}}_{w}$.
Our second goal in this paper is to explain some conjectures related to another interesting variant of the matrix Schubert construction, where $\mathsf{Mat}_{n\times n}$ is replaced by the subset of symmetric matrices $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. Using the same action as in the skew-symmetric case, one may consider the $\mathsf{B}_{n}$-orbits $\mathsf{MSC}^{\mathrm{sym}}_{w}$ in $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ and their closures $\mathsf{MSV}^{\mathrm{sym}}_{w}$, which we call symmetric matrix Schubert varieties. These may again be defined by the rank conditions in (1.1), but restricted to symmetric matrices. When $\mathbb{K}$ does not have characteristic two (as we assume in Example 1.2 below), the corresponding orbit representatives $w$ are provided by all symmetric $n\times n$ partial permutation matrices.
Let $\mathcal{U}^{\mathrm{sym}}=[u_{ij}]_{i,j\in[n]}$ be a symmetric matrix of indeterminates with $u_{ij}=u_{ji}$ and write $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]=\mathbb{K}[u_{ij}:1\leq j\leq i\leq n]$ for the coordinate ring of $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. Then, for any symmetric matrix $w$, we can define an ideal $I^{\mathrm{sym}}_{w}$ of $\mathbb{K}[\mathsf{Mat}_{n\times n}]$ generated by all minors of size $\operatorname{rank}w_{[i][j]}+1$ in $\mathcal{U}_{[i][j]}$ for $i,j\in[n]$. The affine variety $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is the zero set of $I^{\mathrm{sym}}_{w}$.
Now matters become less pleasant: the intersection of two varieties $\mathsf{MSV}^{\mathrm{sym}}_{w}$ need not be a reduced union of varieties $\mathsf{MSV}^{\mathrm{sym}}_{v}$, and no transition recurrence exactly like (1.2) holds with the family of ideals $\textsf{Sym}_{w}$. The next example demonstrates the sort of bad behavior that can occur.
Example 1.2.
The permutation $w=132$ indexes the symmetric matrix Schubert variety
$\mathsf{MSV}^{\mathrm{sym}}_{132}=\{A\in\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}:\operatorname{rank}A_{[2][2]}\leq 1\},\quad\text{which has prime ideal }I^{\mathrm{sym}}_{132}=\langle u_{21}^{2}-u_{11}u_{22}\rangle,$
while $\mathsf{MSV}^{\mathrm{sym}}_{213}=\{A\in\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}:A_{11}=0\}$ has ideal $I^{\mathrm{sym}}_{213}=\langle u_{11}\rangle$. As a variety, $\mathsf{MSV}^{\mathrm{sym}}_{132}\cap\mathsf{MSV}^{\mathrm{sym}}_{213}$ is just the set of symmetric matrices $\left[\begin{smallmatrix}0&0&\ast\\ 0&\ast&\ast\\ \ast&\ast&\ast\end{smallmatrix}\right]$. However, the scheme structure is non-reduced as
$I^{\mathrm{sym}}_{132}+I^{\mathrm{sym}}_{213}=\langle u_{21}^{2}-u_{11}u_{22}\rangle+\langle u_{11}\rangle=\langle u_{11},u_{21}^{2}\rangle.$
Because of obstructions like this one, the set of ideals $I^{\mathrm{sym}}_{w}$ is too small to be assembled into a transition system as defined in Section 3. Nevertheless, we conjecture that a larger transition system exists which includes these ideals. We explain this conjecture precisely in Section 6 (see Conjecture 6.3) and present some evidence that it holds.
There, we also outline some appealing consequences of such a result, for which we have more computational support. In detail, our main conjecture would imply that each ideal $I^{\mathrm{sym}}_{w}$ is both radical and prime (Conjecture 6.7), and that the ideal’s original generating set of minors is a Gröbner basis under an appropriate term order (Conjecture 6.6). These properties have been checked by computer when $\operatorname{\mathsf{char}}(\mathbb{K})=0$ for $n\leq 7$. Conjecture 6.3 would also imply that $I^{\mathrm{sym}}_{w}$ is the ideal of the variety $\mathsf{MSV}^{\mathrm{sym}}_{w}$ (Conjecture 6.5).
In Section 7 we study a more involved application of this last conjectural property. Specifically, assuming Conjecture 6.5 as a hypothesis, we show that the orthogonal Grothendieck polynomials introduced in [28] have well-defined stable limits. This reduces the open problem [28, Prob. 5.3] to the task of constructing a transition system that includes all of the ideals $I^{\mathrm{sym}}_{w}$.
Acknowledgments
This work was partially supported by Hong Kong RGC grants 16306120 and 16304122. We thank Darij Grinberg for helpful comments regarding Lemma 2.9, and Allen Knutson for helping to clarify our approach in §7.
2 Preliminaries
Throughout this article, $\mathbb{K}$ denotes an arbitrary field that is algebraically closed.
2.1 Gröbner bases
Fix a positive integer $N$, let $x_{1},x_{2},\dots,x_{N}$ be commuting variables, and consider the polynomial ring $\mathbb{K}[{\bf x}]:=\mathbb{K}[x_{1},x_{2},\dots,x_{N}]$.
A monomial in $\mathbb{K}[{\bf x}]$ is an element of the form $x_{i_{1}}^{e_{1}}x_{i_{2}}^{e_{2}}\dots x_{i_{l}}^{e_{l}}$ where the indices $i_{1},i_{2},\dots,i_{l}$ are distinct and the exponents $e_{1},e_{2},\dots,e_{l}\geq 0$ are nonnegative. A monomial ideal of $\mathbb{K}[{\bf x}]$ is an ideal generated by a set of monomials. A homogeneous element of $\mathbb{K}[{\bf x}]$ is a $\mathbb{K}$-linear combination of monomials of the same degree. A homogeneous ideal of $\mathbb{K}[{\bf x}]$ is an ideal generated by a set of homogenous elements.
A term order on $\mathbb{K}[{\bf x}]$ is a total order on the set of all monomials, such that $1$ is the unique minimum and such that if $\mathsf{mon}_{1}$, $\mathsf{mon}_{2}$, and $\mathsf{mon}_{3}$ are monomials with $\mathsf{mon}_{1}\leq\mathsf{mon}_{2}$, then $\mathsf{mon}_{1}\mathsf{mon}_{3}\leq\mathsf{mon}_{2}\mathsf{mon}_{3}$. Some common examples of term orders are discussed below.
Example 2.1.
The lexicographic term order on $\mathbb{K}[{\bf x}]$ has $x_{1}^{a_{1}}x_{2}^{a_{2}}\cdots x_{N}^{a_{N}}\leq x_{1}^{b_{1}}x_{2}^{b_{2}}\cdots x_{N}^{b_{N}}$ if and only if $(a_{1},a_{2},\ldots,a_{N})\leq(b_{1},b_{2},\ldots,b_{N})$ in lexicographic order, meaning that if there is a smallest index $i$ with $a_{i}\neq b_{i}$ then $a_{i}<b_{i}$.
Example 2.2.
The (graded) reverse lexicographic term order on $\mathbb{K}[{\bf x}]$ has
$x_{1}^{a_{1}}x_{2}^{a_{2}}\cdots x_{N}^{a_{N}}\leq x_{1}^{b_{1}}x_{2}^{b_{2}}\cdots x_{N}^{b_{N}}$ (2.1)
if and only if both $\sum_{i=1}^{N}a_{i}\leq\sum_{i=1}^{N}b_{i}$ and $(a_{N},\ldots,a_{2},a_{1})\geq(b_{N},\ldots,b_{2},b_{1})$ in lexicographic order.
Remark 2.3.
[…]
  • (a)
    For monomials of different degrees one has $\mathsf{mon}_{1}<\mathsf{mon}_{2}$ if $\deg(\mathsf{mon}_{1})<\deg(\mathsf{mon}_{2})$.
  • (b)
    If $\mathsf{mon}_{1}=\prod_{(i,j)}x_{ij}^{a_{ij}}$ and $\mathsf{mon}_{2}=\prod_{(i,j)}x_{ij}^{b_{ij}}$ are distinct with the same degree then one has $\mathsf{mon}_{1}<\mathsf{mon}_{2}$ when $a_{ij}>b_{ij}$ for the lexicographically largest index $(i,j)$ with $a_{ij}\neq b_{ij}$.
Under this term order, if $\mathsf{mon}_{1}$ and $\mathsf{mon}_{2}$ are both square-free and of the same degree, then $\mathsf{mon}_{1}<\mathsf{mon}_{2}$ if and only there is some variable $x_{ij}$ that does not divide both monomials, and the lexicographically largest such variable divides $\mathsf{mon}_{1}$ but not $\mathsf{mon}_{2}$.
Example 2.4.
Under the reverse lexicographic term order on $\mathbb{K}[x_{11},x_{12},x_{21},x_{22}]$ one has
$1$ $<x_{22}<x_{21}<x_{12}<x_{11}$
$<x_{22}^{2}<x_{21}x_{22}<x_{12}x_{22}<x_{11}x_{22}$
$<x_{21}^{2}<x_{12}x_{21}<x_{11}x_{21}$
$<x_{12}^{2}<x_{11}x_{12}$
$<x_{11}^{2}<\cdots.$
Fix a term order and suppose $f=\sum_{\mathsf{mon}}c_{\mathsf{mon}}\cdot\mathsf{mon}\in\mathbb{K}[{\bf x}]$ where the sum is over all monomials $\mathsf{mon}\in\mathbb{K}[{\bf x}]$ and all but finitely many of the coefficients $c_{\mathsf{mon}}\in\mathbb{K}$ are zero. The initial term $\operatorname{init}(f)$ of $f$ is either the maximal monomial $\mathsf{mon}$ such that $c_{\mathsf{mon}}\neq 0$, or zero when $f=0$.
Choose an ideal $I\subseteq\mathbb{K}[{\bf x}]$. The initial ideal of $I$ is $\operatorname{init}(I):=\mathbb{K}\textnormal{-span}\{\operatorname{init}(f):f\in I\}.$ This vector space is a monomial ideal, and it is a straightforward exercise to check that if $I$ is already a monomial ideal then $\operatorname{init}(I)=I$. A Gröbner basis $G$ for $I$, relative to the chosen term order, is a generating set for $I$ such that $\{\operatorname{init}(g):g\in G\}$ generates $\operatorname{init}(I)$.
Example 2.5.
If $f\in\mathbb{K}[{\bf x}]$ is any polynomial and $I\subseteq\mathbb{K}[{\bf x}]$ is any ideal then
$\operatorname{init}(\langle f\rangle)=\langle\operatorname{init}(f)\rangle\quad\text{and}\quad\operatorname{init}(fI)=\operatorname{init}(f)\operatorname{init}(I).$ (2.2)
Therefore $G=\{f\}$ is a Gröbner basis for the principal ideal $\langle f\rangle$.
Example 2.6.
If $I\subsetneq\mathbb{K}[{\bf x}]$ is a maximal ideal then $I=\langle x_{1}-c_{1},x_{2}-c_{2},\dots,x_{N}-c_{N}\rangle$ for certain constants $c_{i}\in\mathbb{K}$, and for any term order $\operatorname{init}(I)$ is the irrelevant ideal $\langle x_{1},x_{2},\dots,x_{N}\rangle\subsetneq\mathbb{K}[{\bf x}]$. In this case the ideal’s given generating set $G=\{x_{1}-c_{1},x_{2}-c_{2},\dots,x_{N}-c_{N}\}$ is a Gröbner basis.
Not every generating set for an ideal is a Gröbner basis. The ideal $I=\langle x_{1},x_{2}\rangle\subset\mathbb{K}[x_{1},x_{2}]$ is generated by $\{x_{2}-x_{1},x_{2}\}$, but this is not a Gröbner basis for any term order with $x_{1}<x_{2}$. However, if $I\subseteq\mathbb{K}[{\bf x}]$ is any ideal then $\operatorname{init}(\operatorname{init}(I))=\operatorname{init}(I)$. This means that if $S$ is any set of monomials in $\mathbb{K}[{\bf x}]$ then $S$ is a Gröbner basis for the ideal $\langle S\rangle$.
2.2 More on initial ideals
We continue the setup of the previous section. There are a few additional properties of initial ideals that will be of use later. To begin, we recall a few basic facts derived in [31].
Proposition 2.7 (See [31, Prop. 2.4]).
Suppose $I$ and $J$ are ideals in $\mathbb{K}[{\bf x}]$. Then:
  1. (a)
    It always holds that $\operatorname{init}(I)+\operatorname{init}(J)\subseteq\operatorname{init}(I+J)$ and $\operatorname{init}(I\cap J)\subseteq\operatorname{init}(I)\cap\operatorname{init}(J)$.
  2. (b)
    Assume $I\subseteq J$. Then $\operatorname{init}(I)\subseteq\operatorname{init}(J)$, and if $\operatorname{init}(I)=\operatorname{init}(J)$ then $I=J$.
Corollary 2.8.
Any subset $G$ of an ideal $I$ satisfying $\operatorname{init}(I)\subseteq\langle\operatorname{init}(g):g\in G\rangle$ is a Gröbner basis.
Proof.
In this case $\operatorname{init}(I)\subseteq\langle\operatorname{init}(g):g\in G\rangle\subseteq\operatorname{init}(\langle G\rangle)\subseteq\operatorname{init}(I)$ so all containments must be equality. Applying Proposition 2.7(b) to $\langle G\rangle\subseteq I$ shows that $I=\langle G\rangle$ so $G$ is a Gröbner basis. ∎
The next result is a stronger version of [31, Lem. 2.5].
Lemma 2.9.
Suppose $I$ and $J$ are ideals in $\mathbb{K}[{\bf x}]$ with $\operatorname{init}(I+J)=\operatorname{init}(I)+\operatorname{init}(J)$. Then $\operatorname{init}(I\cap J)=\operatorname{init}(I)\cap\operatorname{init}(J)$.
Proof.
The following proof was explained to us by Darij Grinberg. Define a bridge to be a pair $(f,g)$ where $f\in I$ and $g\in J$ and $\operatorname{init}(f)=\operatorname{init}(g)>\operatorname{init}(f-g)$, where >> denotes our term order on monomials and where $\mathsf{mon}>0$ for all monomials. Note that this requires $f$ and $g$ to have the same leading monomial with the same coefficient, unless one has $f=g=0$. Define the floor of a bridge $(f,g)$ to be $\operatorname{init}(f-g)$.
We claim that if $(f,g)$ is a bridge, then $\operatorname{init}(f)=\operatorname{init}(g)\in\operatorname{init}(I\cap J)$. We argue by induction on the bridge floor $\operatorname{init}(f-g)$. The base case when $\operatorname{init}(f-g)=0$ arises if and only if $f=g$, and then the desired property is clear. Assume $f\neq g$. Since $f-g\in I+J$, we have $\operatorname{init}(f-g)\in\operatorname{init}(I+J)=\operatorname{init}(I)+\operatorname{init}(J)$. As $\operatorname{init}I$ and $\operatorname{init}J$ are monomial ideals, the monomial $\operatorname{init}(f-g)$ must belong to $\operatorname{init}(I)$ or $\operatorname{init}(J)$.
First suppose $\operatorname{init}(f-g)\in\operatorname{init}(I)$. Then $\operatorname{init}(f-g)=\operatorname{init}(i)$ for some $0\neq i\in I$. Define $\lambda\in\mathbb{K}$ to be the leading term of fgf-g divided by the leading term of $i$. Then the pair $(f-\lambda i,g)$ is a bridge with a lower floor than $(f,g)$. By induction, we have $\operatorname{init}(f-\lambda i)=\operatorname{init}(g)\in\operatorname{init}(I\cap J)$. Since $\operatorname{init}(f)=\operatorname{init}(f-\lambda i)$, this proves the claim when $\operatorname{init}(f-g)\in\operatorname{init}(I)$. The argument when $\operatorname{init}(f-g)\in\operatorname{init}(J)$ is similar.
This proves the claim. To deduce the lemma, observe that any monomial in $\operatorname{init}(I)\cap\operatorname{init}(J)$ can be written as $\operatorname{init}(f)=\operatorname{init}(g)$ for some bridge $(f,g)$, so is contained in $\operatorname{init}(I\cap J)$ by the claim. Therefore $\operatorname{init}(I)\cap\operatorname{init}(J)\subseteq\operatorname{init}(I\cap J)$, while the reverse containment holds by Proposition 2.7. ∎
Lemma 2.10 ([31, Lem. 2.6]).
If $J\subseteq I$ are homogeneous ideals in $\mathbb{K}[{\bf x}]$ and $f\in\mathbb{K}[{\bf x}]$ is a non-constant homogeneous polynomial such that $I\cap\langle f\rangle=fI$ and $I+\langle f\rangle=J+\langle f\rangle$, then $I=J$.
The conclusion of this lemma does not hold if the ideals involved are inhomogeneous. This can be seen by considering $f=x_{1}\in\mathbb{K}[x_{1},x_{2}]$ with $I=\langle x_{2}\rangle$ and $J=\langle x_{2}-x_{1}x_{2}^{2}\rangle$. Geometrically, the lemma is giving a situation where the properness of a containment of varieties can be tested by intersecting with a hypersurface. In affine space it can happen that one or both intersections are empty, whereas in projective space a hypersurface must intersect a positive-dimensional variety, so it is unsurprising that the lemma is better behaved for homogeneous ideals.
Lemma 2.11.
Suppose $G$ is an additive abelian group with subgroups $A\subseteq B\subseteq G$ and $C\subseteq G$. If $A+C=B+C$ and $A\cap C=B\cap C$ then $A=B$.
Proof.
Let $b\in B$. As $b\in B+C=A+C$ we have $b=a+c$ for some $a\in A$ and $c\in C$. But then $c=b-a\in B\cap C=A\cap C$ so $c\in A$ whence $b\in A$. ∎
3 Transition systems
This section contains our main general results (Definition 3.1 and Theorem 3.3), concerning an inductive approach to computing initial ideals that we refer to as a transition system.
3.1 Main definition
Throughout, we let $A$ be an affine space over an algebraically closed field $\mathbb{K}$. Let $\mathsf{Index}$ be a finite indexing set, choose variables $x_{i}$ to identify $\mathbb{K}[A]=\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ with a ring of polynomials, and then fix a term order on this ring.
For an ideal $I\subseteq\mathbb{K}[A]$ and a nonzero element $0\neq f\in\mathbb{K}[A]$ let
$(I:\langle f\rangle)=\tfrac{1}{f}(I\cap\langle f\rangle)=\{g\in\mathbb{K}[A]:fg\in I\}.$ (3.1)
This ideal quotient is an ideal satisfying $I\subseteq(I:\langle f\rangle)$, with equality if and only if $I\cap\langle f\rangle=fI$. Since $(I:\langle f\rangle)=\mathbb{K}[A]$ if and only if $f\in I$, it follows that if $I$ is maximal with $f\notin I$ then $I=(I:\langle f\rangle)$.
Definition 3.1.
Choose a set $\mathcal{T}$ of pairs $(I,J)$ where $I$ and $J$ are proper ideals of $\mathbb{K}[A]$ such that:
  1. (T1)
    $J$ is a monomial ideal contained in $\operatorname{init}(I)$ with $J=\operatorname{init}(I)$ if $I$ is maximal.
  2. (T2)
    If $(I,J),(I^{\prime},J^{\prime})\in\mathcal{T}$ have $I=I^{\prime}$ then $J=J^{\prime}$.
The set $\mathcal{T}$ is a transition system if for each $(I,J)\in\mathcal{T}$ where $I$ is neither maximal nor equal to $J$, there exists a non-constant $f\in\mathbb{K}[A]$ with $f\notin I$ and a nonempty finite set $\Phi\subseteq\mathcal{T}$ such that:
  1. (T3)
    It holds that $I+\langle f\rangle\subseteq\bigcap_{(P,Q)\in\Phi}P$ and $J+\langle\operatorname{init}(f)\rangle=\bigcap_{(P,Q)\in\Phi}Q$.
  2. (T4)
    If $I^{\prime}:=(I:\langle f\rangle)$ is not equal to $I$ and $J^{\prime}:=(J:\langle\operatorname{init}(f)\rangle)$ then $(I^{\prime},J^{\prime})\in\mathcal{T}$.
Here is a simple example of this construction.
Example 3.2.
Suppose $A=\mathbb{K}$ so that $\mathbb{K}[A]=\mathbb{K}[x]$. Choose elements $c_{i}\in\mathbb{K}$ and define
[…]
$f=x-c_{m}\quad\text{and}\quad\Phi=\left\{(I_{\{m\}},J_{\{m\}})\right\},$
since $(I_{S}:\langle x-c_{m}\rangle)=I_{S\setminus\{m\}}$ and $(J_{S}:\langle\operatorname{init}(x-c_{m})\rangle)=(\langle x^{|S|}\rangle:\langle x\rangle)=\langle x^{|S|-1}\rangle=J_{S\setminus\{m\}}$.
Before discussing more interesting examples, let us present the main application of transition systems: they serve as an inductive tool for verifying initial ideal computations.
Theorem 3.3.
Suppose $\mathcal{T}$ is a transition system. Then $\operatorname{init}(I)=J$ for all $(I,J)\in\mathcal{T}$.
Proof.
If $I$ is maximal then $J=\operatorname{init}(I)$ by assumption (T1), while if $I=J$ then $I$ is a monomial ideal so $J=I=\operatorname{init}(I)$. Suppose $I$ is not maximal and $I\neq J$. Assume $f\notin I$ and $\varnothing\subsetneq\Phi\subseteq\mathcal{T}$ satisfy conditions (T3) and (T4) in Definition 3.1. Then $I\neq I+\langle f\rangle\subseteq\bigcap_{(P,Q)\in\Phi}P$ so
$\operatorname{init}(I)+\langle\operatorname{init}(f)\rangle\subseteq\operatorname{init}\left(I+\langle f\rangle\right)\subseteq\textstyle\operatorname{init}\left(\bigcap_{(P,Q)\in\Phi}P\right)\subseteq\textstyle\bigcap_{(P,Q)\in\Phi}\operatorname{init}(P).$ (3.2)
Since $\mathbb{K}[A]$ is Noetherian and every $(P,Q)\in\Phi$ has $I\subsetneq I+\langle f\rangle\subseteq P$, we may assume by induction that $\operatorname{init}(P)=Q$ for all $(P,Q)\in\Phi$. As we have $J\subseteq\operatorname{init}(I)$ it follows that
$\bigcap_{(P,Q)\in\Phi}\operatorname{init}(P)=\textstyle\bigcap_{(P,Q)\in\Phi}Q=J+\langle\operatorname{init}(f)\rangle\subseteq\operatorname{init}(I)+\langle\operatorname{init}(f)\rangle.$ (3.3)
Thus, the containments in (3.2) and (3.3) must all be equalities, so
$\operatorname{init}(I)+\langle\operatorname{init}(f)\rangle=\operatorname{init}\left(I+\langle f\rangle\right)=J+\langle\operatorname{init}(f)\rangle.$ (3.4)
It remains to prove that $J=\operatorname{init}(I)$. First suppose $(I:\langle f\rangle)=I$ so that $I\cap\langle f\rangle=fI$. Then
$\operatorname{init}(f)\operatorname{init}(I)=\operatorname{init}(fI)$ $=\operatorname{init}(I\cap\langle f\rangle)=\operatorname{init}(I)\cap\langle\operatorname{init}(f)\rangle,$ (3.5)
using Lemma 2.9 and (3.4) for the last equality. Since $J$ and $\operatorname{init}(I)$ are monomial ideals, and therefore homogeneous, and since $\operatorname{init}(f)$ is a non-constant homogenous polynomial, we can use Lemma 2.10 to deduce that $J=\operatorname{init}(I)$. The hypotheses required by the lemma are (3.4) and (3.5).
Alternatively, let $I^{\prime}=(I:\langle f\rangle)$ and $J^{\prime}=(J:\langle\operatorname{init}(f)\rangle)$ and suppose that $I^{\prime}\neq I$. Then we have $(I^{\prime},J^{\prime})\in\mathcal{T}$ by hypothesis and $I\subsetneq I^{\prime}$, so we may assume by induction that $\operatorname{init}(I^{\prime})=J^{\prime}$. Therefore
$\operatorname{init}(I)\cap\langle\operatorname{init}(f)\rangle$ $=\operatorname{init}(I\cap\langle f\rangle)$ by Lemma 2.9 using (3.4)
$=\operatorname{init}(fI^{\prime})$ since $I\cap\langle f\rangle=fI^{\prime}$ by definition
$=\operatorname{init}(f)\operatorname{init}(I^{\prime})$
$=\operatorname{init}(f)J^{\prime}$ by induction
$=J\cap\langle\operatorname{init}(f)\rangle$ by hypothesis.
Given this equation and (3.4), we deduce that $J=\operatorname{init}(I)$ from Lemma 2.11. ∎
The proof just given demonstrates that when $\mathcal{T}$ is a transition system, stronger versions of properties (T3) and (T4) in Definition 3.1 hold. To be precise:
Corollary 3.4.
Suppose $(I,J)$ belongs to a transition system $\mathcal{T}$. Assume $f\notin I$ is a non-constant polynomial and $\Phi\subseteq\mathcal{T}$ is a nonempty finite set satisfying (T3) and (T4) in Definition 3.1. Then:
  1. (a)
    It holds that $I+\langle f\rangle=\bigcap_{(P,Q)\in\Phi}P$ so the containment in (T3) is equality.
  2. (b)
    If $(I:\langle f\rangle)=I$ then $(J:\langle\operatorname{init}(f)\rangle)=J$ so the containment in (T4) is unconditional.
Proof.
Since the containments in (3.2) are all equalities, the two sides of the identity in part (a) have the same initial ideal, so are equal by Proposition 2.7(b). For part (b), note that if $(I:\langle f\rangle)=I$, then (3.5) implies that $(J:\langle\operatorname{init}(f)\rangle)=(\operatorname{init}(I):\langle\operatorname{init}(f)\rangle)=\operatorname{init}(I)=J$. ∎
Another immediate corollary of Theorem 3.3 is the following.
Corollary 3.5.
For a given set $\mathscr{I}$ of ideals in $\mathbb{K}[A]$, there exists at most one transition system $\mathcal{T}$ with $\mathscr{I}=\{I:(I,J)\in\mathcal{T}\}$.
We present one other corollary that relates transition systems to Gröbner bases.
Corollary 3.6.
Suppose $\mathcal{T}$ is a transition system and for each $(I,J)\in\mathcal{T}$ we choose a set $G_{I}\subseteq I$ with $J\subseteq\langle\operatorname{init}(g):g\in G_{I}\rangle$. Then $G_{I}$ is a Gröbner basis for $I$ for each $(I,J)\in\mathcal{T}$.
Proof.
If $(I,J)\in\mathcal{T}$ then $\operatorname{init}(I)=J$ by Theorem 3.3 so $G_{I}$ is a Gröbner basis by Corollary 2.8. ∎
3.2 Examples
Below are some other concrete instances of transition systems. These examples will be considered in greater generality in Sections 4, 5, and 6. The cases here are small enough that one can work out all of the relevant details by hand.
In these examples, $A$ will be some affine space of matrices over $\mathbb{K}$. Given $w\in A$, we let
$\mathsf{MSV}_{w}^{A}=\{M\in A:\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}\text{ for all }i,j\},$
[…]
Example 3.7.
Our first example, while particularly simple, helps illustrate the general setup. Suppose $A$ is the variety of skew-symmetric $3\times 3$ matrices
$A=\left\{\begin{bmatrix}0&-a&-b\\ a&0&-c\\ b&c&0\end{bmatrix}:a,b,c\in\mathbb{K}\right\}.$
Then we may identify $\mathbb{K}[A]=\mathbb{K}[u_{21},u_{31},u_{32}]$ where $u_{ij}(M)=M_{ij}=-M_{ji}$. No skew-symmetric $3\times 3$ matrix is invertible, so the only possible rank tables for elements of $A$ are
$\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\begin{bmatrix}0&0&0\\ 0&0&1\\ 0&1&2\end{bmatrix},\quad\begin{bmatrix}0&0&1\\ 0&0&1\\ 1&1&2\end{bmatrix},\quad\text{and}\quad\begin{bmatrix}0&1&1\\ 1&2&2\\ 1&2&2\end{bmatrix}$.
These rank tables arise from various elements, but in particular from the monomial matrices
$w_{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad w_{2}=\begin{bmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{bmatrix},\quad w_{3}=\begin{bmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{bmatrix},\quad\text{and}\quad w_{4}=\begin{bmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}$.
Define $X_{k}=\mathsf{MSV}^{A}_{w_{k}}\subseteq A$ and $I_{k}=I(X_{k})\subseteq\mathbb{K}[A]$.
We wish to pair each $I_{k}$ with a monomial ideal $J_{k}$ to form a transition system. This is easily done, as $I_{k}$ is already a monomial ideal. To see this, note that we can informally write
$X_{1}=\left\{\begin{bmatrix}0&\\ 0&0\\ 0&0&0\end{bmatrix}\right\},\quad X_{2}=\left\{\begin{bmatrix}0&\\ 0&0\\ 0&\ast&0\end{bmatrix}\right\},\quad X_{3}=\left\{\begin{bmatrix}0&\\ 0&0\\ \ast&\ast&0\end{bmatrix}\right\},\quad\text{and}\quad X_{4}=\left\{\begin{bmatrix}0&\\ \ast&0\\ \ast&\ast&0\end{bmatrix}\right\}$,
so $I_{1}=\langle u_{21},u_{31},u_{32}\rangle\supsetneq I_{2}=\langle u_{21},u_{31}\rangle\supsetneq I_{3}=\langle u_{21}\rangle\supsetneq I_{4}=0$. Then $\mathcal{T}=\{(I_{k},I_{k}):k=1,2,3,4\}$ is transition system relative to any term order on $\mathbb{K}[A]$ since all pairs $(I,J)\in\mathcal{T}$ have $I=J$.
So far we have only seen transition systems $\mathcal{T}$ for which we can take the set $\Phi$ in Definition 3.1 to be a singleton for all $(I,J)\in\mathcal{T}$. The next example presents a case where this is not possible.
Example 3.8.
Let $A=\left\{\begin{bmatrix}a&b\\ c&d\end{bmatrix}:a,b,c,d\in\mathbb{K}\right\}$ so $\mathbb{K}[A]=\mathbb{K}[x_{11},x_{12},x_{21},x_{22}]$ where $x_{ij}(M)=M_{ij}$. The possible rank tables for elements of $A$ arise from the partial permutation matrices
$w_{1}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix},\quad w_{2}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\quad w_{3}=\begin{bmatrix}0&0\\ 1&0\end{bmatrix},\quad w_{4}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix},\quad w_{5}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad w_{6}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad w_{7}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$.
Define $X_{k}=\mathsf{MSV}_{w_{k}}^{A}\subseteq A$ and $I_{k}=I(X_{k})\subseteq\mathbb{K}[A]$. Choose any term order on $\mathbb{K}[A]$. Then let
$J_{k}=\operatorname{init}(I_{k})\quad\text{and}\quad\mathcal{T}=\{(I_{k},J_{k}):k=1,2,3,4,5,6,7\}.$
We now investigate whether $\mathcal{T}$ is a transition system. As one can informally express
$X_{1}$ $=\left\{\begin{bmatrix}0&0\\ 0&0\end{bmatrix}\right\},\quad X_{2}=\left\{\begin{bmatrix}0&0\\ 0&\ast\end{bmatrix}\right\},\quad X_{3}=\left\{\begin{bmatrix}0&0\\ \ast&\ast\end{bmatrix}\right\},\quad X_{4}=\left\{\begin{bmatrix}0&\ast\\ 0&\ast\end{bmatrix}\right\}$,
$X_{5}$ $=\left\{\begin{bmatrix}0&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{6}=\left\{\begin{bmatrix}\ast&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{7}=\left\{\det=0\right\},$
it follows that
$I_{1}=\langle x_{11},x_{12},x_{21},x_{22}\rangle,$ $I_{2}=\langle x_{11},x_{12},x_{21}\rangle,\quad$ $I_{3}=\langle x_{11},x_{12}\rangle,$ $I_{4}=\langle x_{11},x_{21}\rangle,$
$I_{5}=\langle x_{11}\rangle$ $I_{6}=0,$ $I_{7}=\langle x_{11}x_{22}-x_{12}x_{21}\rangle.$
Therefore $J_{k}=\operatorname{init}(J_{k})=I_{k}$ for $k\in[6]$ while $J_{7}=\operatorname{init}(I_{7})$ is either $\langle x_{12}x_{21}\rangle$ or $\langle x_{11}x_{22}\rangle$, depending on the choice of term order.
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{7},J_{7})$. First suppose $J_{7}=\langle x_{12}x_{21}\rangle$. Then
$I_{7}+\langle x_{11}\rangle=J_{7}+\langle x_{11}\rangle=\langle x_{11},x_{12}x_{21}\rangle=I_{3}\cap I_{4}=J_{3}\cap J_{4},$
so $f=x_{11}$ and $\Phi=\{(I_{3},J_{3}),(I_{4},J_{4})\}$ satisfy condition (T3) in Definition 3.1. For these choices, condition (T4) holds vacuously, and we conclude that $\mathcal{T}$ is a transition system.
For term orders with $J_{7}=\langle x_{11}x_{22}\rangle$, the set $\mathcal{T}$ is not a transition system, but it can be extended to one by adding the ideal pair $(I_{8},J_{8})$ with $I_{8}=J_{8}=\langle x_{12},x_{22}\rangle$. Then, when $(I,J)=(I_{7},J_{7})$, conditions (T3) and (T4) are satisfied by taking $f=x_{12}$ and $\Phi=\{(I_{3},J_{3}),(I_{8},J_{8})\}$. The ideal $I_{8}$ is radical, but it corresponds to a closed subvariety of $A$ not of the form $\mathsf{MSV}_{w}^{A}$.
The next example shows a case where we cannot form a transition system including $I(\mathsf{MSV}_{w}^{A})$ for all $w\in A$ without adding non-radical ideals.
Example 3.9.
Let $A=\left\{\begin{bmatrix}a&b\\ b&c\end{bmatrix}:a,b,c\in\mathbb{K}\right\}$ by the variety of symmetric $2\times 2$ matrices over $\mathbb{K}$ so that $\mathbb{K}[A]=\mathbb{K}[u_{11},u_{21},u_{22}]$ with $u_{ij}(M)=M_{ij}=M_{ji}$. We will just consider $\mathbb{K}[A]$ under the reverse lexicographic term order. The possible rank tables for elements of $A$ arise from the matrices
$w_{1}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},\quad w_{2}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\quad w_{3}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad w_{4}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\quad w_{5}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$.
Let $X_{k}=\mathsf{MSV}_{w_{k}}^{A}\subseteq A$ and $I_{k}=I(X_{k})\subseteq\mathbb{K}[A]$. We can informally write the subvarieties $X_{k}$ as
$X_{1}=\left\{\det=0\right\},\quad X_{2}=\left\{\begin{bmatrix}\ast&\\ \ast&\ast\end{bmatrix}\right\},\quad X_{3}=\left\{\begin{bmatrix}0&\\ \ast&\ast\end{bmatrix}\right\},\quad X_{4}=\left\{\begin{bmatrix}0&\\ 0&\ast\end{bmatrix}\right\},\quad X_{5}=\left\{\begin{bmatrix}0&\\ 0&0\end{bmatrix}\right\}$,
and so $I_{1}=\langle u_{21}^{2}-u_{11}u_{22}\rangle$ and $I_{2}=0\subsetneq I_{3}=\langle u_{11}\rangle\subsetneq I_{4}=\langle u_{11},u_{21}\rangle\subsetneq I_{5}=\langle u_{11},u_{21},u_{22}\rangle$. Define
$J_{1}=\operatorname{init}(I_{1})=\langle u_{21}^{2}\rangle\quad\text{and}\quad J_{k}=\operatorname{init}(I_{k})=I_{k}\text{ for $k\in\{2,3,4,5\}$}.$
The set $\{(I_{k},J_{k}):k=1,2,3,4,5\}$ fails to be a transition system, though only by a small margin.
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{1},J_{1})$ since $I_{k}=J_{k}$ when $k>1$. The problem is that $J_{1}+\langle\operatorname{init}(f)\rangle$ is never equal to $\bigcap_{k\in S}J_{k}$ for any choice of set $\varnothing\subsetneq S\subseteq\{1,2,3,4,5\}$ when $f\in\mathbb{K}[A]$ is non-constant with $f\notin I_{1}$. To get around this, define
$I_{6}=J_{6}=\langle u_{11},u_{21}^{2}\rangle=I_{1}+\langle u_{11}\rangle=J_{1}+\langle u_{11}\rangle.$
Now if $(I,J)=(I_{1},J_{1})$ then (T3) holds for $f=u_{11}\notin I_{1}$ and $\Phi=\{(I_{6},J_{6})\}$, and $(I:\langle f\rangle)=I$ so (T4) holds vacuously. Thus the set $\mathcal{T}=\{(I_{k},J_{k}):k\in[6]\}$ is a transition system. Notice, however, that the extra ideal $I_{6}$ is not radical, so corresponds to a non-reduced subscheme of $A$.
3.3 Orbit closures
We are most interested in transition systems $\mathcal{T}$ made up of pairs $(I,J)$ where $I$ ranges over the orbit closures of certain maximal ideals. This section studies the relevant orbit closure operation. It is convenient to present these results in an entirely self-contained way, but generalizations of this material in terms of equivariant sheaves may be well-known; see, e.g., [5, 32]. If the reader is willing to believe that obvious properties of orbit closures of subvarieties extend naturally from radical ideals to all ideals, the details of this section can be skipped.
Continue to let $A$ be an affine space over an algebraically closed field $\mathbb{K}$. Suppose $G$ is a connected linear algebraic group over $\mathbb{K}$ that acts algebraically on $A$. We identify the coordinate ring $\mathbb{K}[G\times A]$ with $\mathbb{K}[G]\otimes\mathbb{K}[A]$ where $\otimes$ denotes the tensor product over $\mathbb{K}$. The action of $G$ on $A$ determines a map
$\begin{array}[]{rrcl}\alpha:&G\times A&\to&A\\ &(g,a)&\mapsto&g\cdot a\end{array}\quad\text{whose pullback is }\quad\begin{array}[]{rrcl}\alpha^{*}:&\mathbb{K}[A]&\to&\mathbb{K}[G\times A]\\ &f&\mapsto&f\circ\alpha.\end{array}$ (3.6)
Since the action of $G$ on $A$ is algebraic, $\alpha$ is a morphism of algebraic varieties and $\alpha^{*}$ is a ring homomorphism. In addition, $\alpha^{\ast}$ is surjective as $\alpha^{*}(f)$ sends $(1,a)\mapsto f(a)$ for all $a\in A$.
When $I\subseteq\mathbb{K}[A]$ is an ideal, we denote the ideal of $\mathbb{K}[G\times A]$ generated by the set $\{1\otimes f:f\in I\}$ as $\mathbb{K}[G]\otimes I$. Since $\mathbb{K}[G]$ is a free and flat $\mathbb{K}$-module, the following holds:
Lemma 3.10.
If $I$ is any ideal in $\mathbb{K}[A]$ then $\mathbb{K}[G\times A]/(\mathbb{K}[G]\otimes I)\cong\mathbb{K}[G]\otimes(\mathbb{K}[A]/I)$.
Now, the orbit closure of an ideal $I\subseteq\mathbb{K}[A]$, relative to the action of $G$, is defined to be
$\mathrm{cl}_{G}(I)=(\alpha^{*})^{-1}(\mathbb{K}[G]\otimes I)\subseteq\mathbb{K}[A].$ (3.7)
Because $\alpha^{*}$ is a ring homomorphism, this preimage is an ideal in $\mathbb{K}[A]$.
Example 3.11.
If $I=0$ then $\mathrm{cl}_{G}(0)=0$ and if $I=\mathbb{K}[A]$ then $\mathrm{cl}_{G}(I)=\mathbb{K}[A]$.
For any ideal $I\subseteq\mathbb{K}[A]$ define $V(I):=\{a\in A:f(a)=0\text{ for all }f\in I\}$ and for any subset $V\subseteq A$ define $I(V):=\{f\in\mathbb{K}[A]:f(v)=0\text{ for all }v\in V\}.$ Because $\mathbb{K}$ is algebraically closed, Hilbert’s Nullstellensatz asserts that if $I$ is any ideal then $V(I)$ is a closed affine subvariety and $I(V(I))=\sqrt{I}$ is equal to the radical of $I$.
The ideal $\mathrm{cl}_{G}(I)$ is related to the closure of the union of the $G$-orbits through all points in $V(I)$.
Proposition 3.12.
If $I$ is a radical ideal in $\mathbb{K}[A]$ then so is $\mathrm{cl}_{G}(I)$, and in this case $V(\mathrm{cl}_{G}(I))$ is the closure of the set $\{g\cdot a:(g,a)\in G\times V(I)\}$ in the Zariski topology of $A$.
Proof.
An ideal $I$ in a commutative ring $R$ is radical if and only if the quotient ring $R/I$ is reduced in the sense of having no nonzero nilpotent elements. Since a linear algebraic group is an affine variety, the coordinate ring $\mathbb{K}[G]$ is an integral domain and therefore reduced. On the other hand, by Lemma 3.10 we know that $\mathbb{K}[G\times A]/(\mathbb{K}[G]\otimes I)\cong\mathbb{K}[G]\otimes(\mathbb{K}[A]/I)$. Tensor products of reduced algebras over perfect fields are always reduced [3, Thm. 3, Chapter V, §15], and our field $\mathbb{K}$ is perfect since it is algebraically closed. Thus if $I$ is radical then both $\mathbb{K}[A]/I$ and $\mathbb{K}[G]\otimes(\mathbb{K}[A]/I)$ are reduced, so $\mathbb{K}[G]\otimes I$ is radical, as is its preimage $\mathrm{cl}_{G}(I)$ under the ring homomorphism $\alpha^{*}$.
Now let $\mathcal{O}=\{g\cdot a:(g,a)\in G\times V(I)\}\subseteq A$. If $f\in\mathrm{cl}_{G}(I)$ then $\alpha^{*}(f)\in\mathbb{K}[G]\otimes I$, so for all $g\in G$ and $a\in V(I)$ we have $f(g\cdot a)=\alpha^{*}(f)(g,a)=0$. This shows that $\mathrm{cl}_{G}(I)\subseteq V(\mathcal{O})$ so $\mathcal{O}\subseteq V(\mathrm{cl}_{G}(I))$. Since $V(\mathrm{cl}_{G}(I))$ is a closed affine subvariety, it also contains the closure of $\mathcal{O}$.
Conversely, if $J\subseteq\mathbb{K}[A]$ is any radical ideal with $\mathcal{O}\subseteq V(J)$, then each $f\in J$ has
$\alpha^{*}(f)\in I(G\times V(I))=V(G)\otimes\mathbb{K}[A]+\mathbb{K}[G]\otimes I(V(I))=\mathbb{K}[G]\otimes I.$
Thus $J\subseteq\mathrm{cl}_{G}(I)$ and $V(\mathrm{cl}_{G}(I))\subseteq V(J)$, so the closure of $\mathcal{O}$, which is the intersection of all such sets $V(J)$, also contains $V(\mathrm{cl}_{G}(I))$. ∎
The reduction of a commutative ring $R$ is the quotient $R^{\prime}:=R/\sqrt{0}$ by its nilradical. An ideal $I$ in $R$ is primary if all zero divisors in $R/I$ are nilpotent, or equivalently if $(R/I)^{\prime}$ is an integral domain. The closure operation $\mathrm{cl}_{G}$ preserves this property.
Proposition 3.13.
If $I$ is a primary ideal in $\mathbb{K}[A]$ then so is $\mathrm{cl}_{G}(I)$.
Proof.
If $M$ and $N$ are rings that are commutative $\mathbb{K}$-algebras, then $(M\otimes N)^{\prime}\cong(M^{\prime}\otimes N^{\prime})^{\prime}$. Hence, for any ideal $I\subseteq\mathbb{K}[A]$, Lemma 3.10 implies that the reduction of $\mathbb{K}[G\times A]/(\mathbb{K}[G]\otimes I)$ is isomorphic to the reduction of $\mathbb{K}[G]^{\prime}\otimes(\mathbb{K}[A]/I)^{\prime}$.
Since $G$ is connected, it holds that $\mathbb{K}[G]^{\prime}=\mathbb{K}[G]$ is an integral domain, and if $I$ is primary then $(\mathbb{K}[A]/I)^{\prime}$ is an integral domain. In this case $\mathbb{K}[G]^{\prime}\otimes(\mathbb{K}[A]/I)^{\prime}$ is a tensor product of integral domains over an algebraically closed field, so is itself an integral domain that is equal to its reduction. Thus if $I$ is primary then so is $\mathbb{K}[G]\otimes I$, as is the preimage $\mathrm{cl}_{G}(I)$. ∎
We note some more general properties of the orbit closure operation $\mathrm{cl}_{G}$.
Lemma 3.14.
If $I\subseteq\mathbb{K}[A]$ is any ideal then $\mathrm{cl}_{G}(I)\subseteq I$.
Proof.
Let $\iota:A\to G\times A$ be the map $a\mapsto(1,a)$. Then $\iota^{*}(f_{1}\otimes f_{2})=(f_{1}\otimes f_{2})\circ\iota=f_{1}(1)f_{2}$ when $f_{1}\in\mathbb{K}[G]$ and $f_{2}\in\mathbb{K}[A],$ so $\iota^{*}$ is a surjection $\mathbb{K}[G]\otimes I\to I$. Hence, if $f\in\mathrm{cl}_{G}(I)$ then $\iota^{*}\circ\alpha^{*}(f)\in\iota^{*}(\mathbb{K}[G]\otimes I)=I.$ But $\iota^{*}\circ\alpha^{*}=(\alpha\circ\iota)^{*}=\mathrm{id}_{\mathbb{K}[A]}$, so if $f\in\mathrm{cl}_{G}(I)$ then $f\in I$. ∎
Lemma 3.15.
If $I\subseteq\mathbb{K}[A]$ is any ideal then $\mathrm{cl}_{G}(\mathrm{cl}_{G}(I))=\mathrm{cl}_{G}(I)$.
Proof.
Let $\pi:G\times A\to A$ be the projection map onto the second factor $(g,a)\mapsto a$. Then $\pi^{*}(f)=1\otimes f$ so $\pi^{*}\circ\iota^{*}(f_{1}\otimes f_{2})=f_{1}(1)\otimes f_{2}\in\mathbb{K}[G]\otimes\mathbb{K}[A]$ for $f_{1}\in\mathbb{K}[G]$ and $f_{2}\in\mathbb{K}[A].$ Thus $\pi^{*}\circ\iota^{*}$ evaluates the first factor of an element of $\mathbb{K}[G]\otimes\mathbb{K}[A]$ at $1\in G$, so this composition restricts for any ideal $I\subseteq\mathbb{K}[A]$ to a surjective $\mathbb{K}$-linear map $\mathbb{K}[G]\otimes I\to\mathbb{K}\otimes I$.
Since $\mathrm{cl}_{G}(I)$ is defined to be $(\alpha^{*})^{-1}\langle\pi^{*}(I)\rangle$ where $\langle S\rangle$ is the ideal generated by $S$, we have $\mathrm{cl}_{G}(\mathrm{cl}_{G}(I))=(\alpha^{*})^{-1}\left\langle\pi^{*}((\alpha^{*})^{-1}(\mathbb{K}[G]\otimes I))\right\rangle.$ To simplify this expression, we now prove that
$\pi^{*}((\alpha^{*})^{-1}(f))=\{\pi^{*}\circ\iota^{*}(f)\}\quad\text{for any $f\in\mathbb{K}[G]\otimes\mathbb{K}[A]$.}$
Indeed, suppose $F\in\pi^{*}((\alpha^{*})^{-1}(f))$. Then $F=\pi^{*}(p)=1\otimes p$ for some $p\in\mathbb{K}[A]$ with $\alpha^{*}(p)=f$. This means that if $g\in G$ and $a\in A$ then $p(ga)=f(g,a)$, so
$F(g,a)=(1\otimes p)(g,a)=p(a)=f(1,a)=f\circ\iota\circ\pi(g,a).$
This shows that $F=\pi^{*}\circ\iota^{*}(f)$. Using our claim, we deduce that
$\mathrm{cl}_{G}(\mathrm{cl}_{G}(I))=(\alpha^{*})^{-1}\left\langle\pi^{*}\circ\iota^{*}(\mathbb{K}[G]\otimes I)\right\rangle=(\alpha^{*})^{-1}(\langle\mathbb{K}\otimes I\rangle)=(\alpha^{*})^{-1}(\mathbb{K}[G]\otimes I)=\mathrm{cl}_{G}(I).$
We define an ideal $I\subseteq\mathbb{K}[A]$ to be $G$-stable if $\mathrm{cl}_{G}(I)=I$. The previous lemma implies that $\mathrm{cl}_{G}(I)$ is $G$-stable and contained in $I$. It turns out that $\mathrm{cl}_{G}(I)$ is the largest ideal of this type.
Proposition 3.16.
If $I\subseteq\mathbb{K}[A]$ is any ideal then $\mathrm{cl}_{G}(I)$ is the sum of all $G$-stable ideals in $I$.
Proof.
Let $\Sigma$ be the sum of all $G$-stable ideals contained in $I$. Since $\mathrm{cl}_{G}(I)$ is $G$-stable and contained in $I$ by Lemmas 3.14 and 3.15, we have $\mathrm{cl}_{G}(I)\subseteq\Sigma$. Conversely, if $J\subseteq I$ then it is clear that $\mathrm{cl}_{G}(J)\subseteq\mathrm{cl}_{G}(I)$, so if $J$ is also $G$-stable then $J=\mathrm{cl}_{G}(J)\subseteq\mathrm{cl}_{G}(I)$. Hence $\Sigma\subseteq\mathrm{cl}_{G}(I)$. ∎
The group $G$ acts on $\mathbb{K}[A]$ by the formula $g\cdot f:a\mapsto f(g^{-1}\cdot a)$ for $g\in G$, $f\in\mathbb{K}[A]$, and $a\in A$. Our notion of $G$-stability can alternatively be characterized in terms of this action.
Proposition 3.17.
An ideal $I\subseteq\mathbb{K}[A]$ is $G$-stable if and only if $g\cdot I=I$ for all $g\in G$.
Proof.
Fix $g\in G$ and suppose $f\in\mathrm{cl}_{G}(I)$. Then we can write $\alpha^{*}(f)=\sum_{i}p_{i}\otimes q_{i}$ as a finite sum with each $p_{i}\in\mathbb{K}[G]$ and $q_{i}\in I$. The group $G$ also acts on $\mathbb{K}[G]$, and since
$\alpha^{*}(g\cdot f)(x,y)=f(g^{-1}x\cdot y)=\alpha^{*}(f)(g^{-1}x,y)\quad\text{for }x\in G\text{ and }y\in A,$
we have $\alpha^{*}(g\cdot f)=\sum_{i}(g\cdot p_{i})\otimes q_{i}\in\mathbb{K}[G]\otimes I.$ Thus $g\cdot f\in\mathrm{cl}_{G}(I)$, so if $I=\mathrm{cl}_{G}(I)$ then $g\cdot I\subseteq I$. As $G$ is a group, having $g\cdot I\subseteq I$ for all $g\in G$ implies that $g\cdot I=I$ for all $g\in G$.
Now suppose $I$ is an ideal with $g\cdot I=I$ for all $g\in G$. In view of Lemma 3.14 it suffices to show that $I\subseteq\mathrm{cl}_{G}(I)$. Again let $f\in\mathrm{cl}_{G}(I)$. We must check that $\alpha^{*}(f)\in\mathbb{K}[G]\otimes I$. To this end, notice that we can write $\alpha^{*}(f)=\sum_{i=1}^{m}p_{i}\otimes q_{i}$ where $p_{1},p_{2},\dots,p_{m}\in\mathbb{K}[G]$ are linearly independent over $\mathbb{K}$ and each $q_{i}\in\mathbb{K}[A]$. For $g\in G$ let $\iota_{g}:A\to G\times A$ denote the map $a\mapsto(g,a)$. Then
$\iota_{g}^{*}\circ\alpha^{*}(f)=\alpha^{*}(f)\circ\iota_{g}=\sum_{i=1}^{m}p_{i}(g)q_{i}\in\mathbb{K}[A].$
At the same time, for each $g\in G$ we have $\iota_{g}^{*}\circ\alpha^{*}(f)=f\circ\alpha\circ\iota_{g}=g^{-1}\cdot f$ which is in $I$ by hypothesis.
Finally, observe that as $p_{1},p_{2},\dots,p_{m}$ are linearly independent as functions on $G$, we may choose group elements $g_{1},g_{2},\dots,g_{m}\in G$ such that the matrix $\left[p_{i}(g_{j})\right]_{1\leq i,j\leq m}$ is invertible. Then the fact that we have $\sum_{i=1}^{m}p_{i}(g_{j})q_{j}=\iota_{g}^{*}\circ\alpha^{*}(f)\in I$ for all $j$ implies that the functions $q_{1},q_{2},\dots,q_{m}$ are all in $I$, and so $\alpha^{*}(f)=\sum_{i=1}^{m}p_{i}\otimes q_{i}\in\mathbb{K}[G]\otimes I$ as needed. ∎
3.4 Reductions
Transition systems that arise from orbit closures of maximal ideals are simpler than the general case in two respects. First, we can usually choose $f\notin I$ in Definition 3.1 to be one of the variables $x_{i}$ generating $\mathbb{K}[A]$. Second, condition (T4) in Definition 3.1 will often hold by default.
We prove two technical lemmas in this section in order to efficiently make these reductions later. Continue to let $G$ be a linear algebraic group that acts algebraically on an affine space $A$, with $G$ and $A$ both defined over any algebraically closed field $\mathbb{K}$. Recall that we identify $\mathbb{K}[A]$ with the polynomial ring $\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ where $\mathsf{Index}$ is some finite indexing set.
Lemma 3.18.
Suppose there exists a partial order $\preceq$ on $\mathsf{Index}$ such that if $j\in\mathsf{Index}$ then
$\alpha^{\ast}(x_{j})=\sum_{i\preceq j}f_{ij}\otimes x_{i}\in\mathbb{K}[G\times A]\quad\text{for elements $f_{ij}\in\mathbb{K}[G]$ with $f_{jj}$ invertible}.$
Let $I$ be an ideal of $\mathbb{K}[A]$ and suppose $j\in\mathsf{Index}$ has the property that $x_{i}\in I$ for all $i\prec j$. Then:
  1. (a)
    We have $x_{j}\notin\mathrm{cl}_{G}(I)$ if and only if $x_{j}\notin I$.
  2. (b)
    It holds that $(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)=\mathrm{cl}_{G}((I:\langle x_{j}\rangle)).$
Proof.
If $x_{j}\notin I$ then $x_{j}\notin\mathrm{cl}_{G}(I)$ since $\mathrm{cl}_{G}(I)\subseteq I$. Conversely, if $x_{j}\in I$ then $x_{i}\in I$ for all $i\preceq j$ so by hypothesis $\alpha^{\ast}(x_{j})\in\mathbb{K}[G]\otimes I$ and $x_{j}\in\mathrm{cl}_{G}(I)$. This proves part (a).
To prove part (b), let $I^{\prime}=(I:\langle x_{j}\rangle)$. Note that if $x_{j}\in I$ then we have $I^{\prime}=\mathbb{K}[A]=\mathrm{cl}_{G}(I^{\prime})$ and also $(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)=\mathbb{K}[A]$ since $x_{j}\in\mathrm{cl}_{G}(I)$ by part (a). We may therefore assume that $x_{j}\notin I$, although this condition does not play a role in the following argument. To show that $(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)=\mathrm{cl}_{G}(I^{\prime})$ we check that each ideal contains the other. First let $F\in\mathrm{cl}_{G}(I^{\prime})$. Then
$\alpha^{\ast}(x_{j}F)=\alpha^{\ast}(x_{j})\alpha^{\ast}(F)=\sum_{i\prec j}\underbrace{(f_{ij}\otimes x_{i})}_{\in\mathbb{K}[G]\otimes I}\alpha^{\ast}(F)+\underbrace{(f_{jj}\otimes x_{j})\alpha^{\ast}(F)}_{\in\mathbb{K}[G]\otimes x_{j}I^{\prime}}.$
As we have $x_{j}I^{\prime}=I\cap\langle x_{j}\rangle\subseteq I$, it follows that $\alpha^{\ast}(x_{j}F)\in\mathbb{K}[G]\otimes I$, so we deduce that $x_{j}F\in\mathrm{cl}_{G}(I)$ and therefore $F\in(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)$. This shows that $\mathrm{cl}_{G}(I^{\prime})\subseteq(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)$.
For the reverse containment, consider a generic element of $\mathrm{cl}_{G}(I)\cap\langle x_{j}\rangle$, which must have the form $x_{j}F$ for some $F\in\mathbb{K}[A]$ with $\alpha^{\ast}(x_{j}F)\in\mathbb{K}[G]\otimes I$. Since we can write
$\alpha^{\ast}(x_{j}F)=\alpha^{\ast}(x_{j})\alpha^{\ast}(F)=\sum_{i\prec j}\underbrace{(f_{ij}\otimes x_{i})}_{\in\mathbb{K}[G]\otimes I}\alpha^{\ast}(F)+(f_{jj}\otimes x_{j})\alpha^{\ast}(F),$
it follows that $(f_{jj}\otimes x_{j})\alpha^{\ast}(F)\in\mathbb{K}[G]\otimes I$. As $f_{jj}$ is invertible, we have $(1\otimes x_{j})\alpha^{\ast}(F)\in\mathbb{K}[G]\otimes I$. This element is also in $\mathbb{K}[G]\otimes\langle x_{j}\rangle$, and so
$(1\otimes x_{j})\alpha^{\ast}(F)\in(\mathbb{K}[G]\otimes I)\cap(\mathbb{K}[G]\otimes\langle x_{j}\rangle).$ (3.8)
Recall that we view both $\mathbb{K}[G]\otimes I$ and $\mathbb{K}[G]\otimes\langle x_{j}\rangle$ as submodules of $\mathbb{K}[G\times A]$, so the intersection on the right is well-defined. Under this convention, it holds that
$(\mathbb{K}[G]\otimes I)\cap(\mathbb{K}[G]\otimes\langle x_{j}\rangle)=\mathbb{K}[G]\otimes(I\cap\langle x_{j}\rangle)=\mathbb{K}[G]\otimes x_{j}I^{\prime}$ (3.9)
since $\mathbb{K}[G]$ is a free $\mathbb{K}$-module (as it is well-known for modules over a commutative ring that tensoring with a flat module commutes with finite intersections). From (3.8) and (3.9) we get
$(1\otimes x_{j})\alpha^{\ast}(F)\in\mathbb{K}[G]\otimes x_{j}I^{\prime},$
which is only possible if $\alpha^{\ast}(F)\in\mathbb{K}[G]\otimes I^{\prime}$. We conclude that $F\in\mathrm{cl}_{G}(I^{\prime})$ and $x_{j}F\in x_{j}\mathrm{cl}_{G}(I^{\prime})$. This means that $\mathrm{cl}_{G}(I)\cap\langle x_{j}\rangle\subseteq x_{j}\mathrm{cl}_{G}(I^{\prime})$ so $(\mathrm{cl}_{G}(I):\langle x_{j}\rangle)\subseteq\mathrm{cl}_{G}(I^{\prime})$. ∎
Our second lemma is a straightforward property of maximal ideals in polynomial rings.
Lemma 3.19.
Suppose $I=\langle x_{i}-c_{i}:i\in\mathsf{Index}\rangle\subset\mathbb{K}[A]$ for certain constants $c_{i}\in\mathbb{K}$ with $i\in\mathsf{Index}$, so that $I$ is the (maximal) vanishing ideal of a point in $A$. Then for any given $j\in\mathsf{Index}$ the following properties are equivalent: (a) $x_{j}\notin I$, (b) $(I:\langle x_{j}\rangle)=I$, and (c) $c_{j}\neq 0$.
Proof.
Properties (a) and (b) are equivalent since $I$ is a maximal ideal, in view of the observations before Definition 3.1. If $x_{j}\notin I$ then clearly $c_{j}\neq 0$. Conversely, if $c_{j}\neq 0$ then we cannot have $x_{j}\in I$ as then $1=\tfrac{1}{c_{j}}(x_{j}-(x_{j}-c_{j}))\in I$, contradicting the assumption that $I$ is a proper ideal. ∎
4 Matrix Schubert varieties
[…]
Definition 4.1.
For each $m\times n$ matrix $w$ over $\mathbb{K}$, the corresponding matrix Schubert cell and matrix Schubert variety are the respective subsets of $\mathsf{Mat}_{m\times n}$ defined by the rank conditions
$\mathsf{MSC}_{w}$ $=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$ (4.1)
$\mathsf{MSV}_{w}$ $=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$
where $M_{[i][j]}$ stands for the upper left $i\times j$ submatrix of $M$.
To study matrix Schubert varieties, one is naturally led to investigate the vanishing ideals $I(\mathsf{MSV}_{w})$. From this perspective, it is of interest to compute the initial ideal of $I(\mathsf{MSV}_{w})$ and then to construct a corresponding Gröbner basis. We explain below how to do such calculations using transition systems. This will recover several results from [25].
The transition system approach is best suited to term orders on $\mathbb{K}[\mathsf{Mat}_{m\times n}]$ with a certain antidiagonal property defined in [25, §1.5]. For simplicity, we will always work with the following instance of such an order. First, identify the coordinate ring of $\mathsf{Mat}_{m\times n}$ with
$\mathbb{K}[\mathsf{Mat}_{m\times n}]=\mathbb{K}[x_{ij}:1\leq i\leq m,1\leq j\leq n]$
by setting $x_{ij}$ to be the linear function on matrices given by $x_{ij}(M)=M_{ij}\in\mathbb{K}$. Then order the monomials in this ring using the reverse lexicographic term order explained in Remark 2.3.
Remark 4.2.
Prior to [25], the problem of finding Gröbner bases for various special cases of the ideals $I(\mathsf{MSV}_{w})$ had already appeared in a number of places; see [25, §2.4] for a historical overview. Since [25], there has only been partial progress on understanding the Gröbner geometry of matrix Schubert varieties for arbitrary term orders; see, for example, [14, 19, 24, 26].
4.1 A reverse lexicographic transition system
The variety $\mathsf{MSV}_{w}$ is irreducible and equal to the Zariski closure of $\mathsf{MSC}_{w}$ [32, Ch. 15]. Let $\mathsf{B}_{m}$ and $\mathsf{B}_{n}$ be the Borel subgroups of invertible lower-triangular matrices in $\mathsf{Mat}_{m\times m}$ and $\mathsf{Mat}_{n\times n}$. Then $\mathsf{B}_{m}\times\mathsf{B}_{n}$ acts algebraically on $\mathsf{Mat}_{m\times n}$ by $(g,h)\cdot M=gMh^{\top}$, and $\mathsf{MSC}_{w}$ is the orbit of $w$ under this action. It follows that in the notation of Section 3.3, we can express
$I(\mathsf{MSV}_{w})=\mathrm{cl}_{\mathsf{B}_{m}\times\mathsf{B}_{n}}(\mathcal{M}_{w})\quad\text{for the maximal ideal }\mathcal{M}_{w}:=\langle x_{ij}-w_{ij}:(i,j)\in[m]\times[n]\rangle.$ (4.2)
As a matrix has rank $k$ if and only if all of its $(k+1)\times(k+1)$ minors are zero, $I(\mathsf{MSV}_{w})$ contains certain minors of the matrix of variables $\mathcal{X}=[x_{ij}]_{1\leq i\leq m,1\leq j\leq n}$. Specifically, if we define
$I_{w}:=\left\langle\det(\mathcal{X}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle$ (4.3)
then it is clear a priori that $I_{w}\subseteq I(\mathsf{MSV}_{w})$. We will see later (in Corollary 4.7) that this containment is actually equality.
For a $k\times k$ matrix $M$ define $\operatorname{adiag}(M):=M_{1,k}M_{2,k-1}M_{3,k-2}\cdots M_{k,1}$ to be the product of the entries on the antidiagonal. Then we also consider the monomial ideal
$J_{w}:=\left\langle\operatorname{adiag}(\mathcal{X}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle.$ (4.4)
Observe that $\operatorname{adiag}(\mathcal{X}_{RC})$ is the leading term of $\det(\mathcal{X}_{RC})$ under the reverse lexicographic term order. Therefore, relative to this term order, we automatically have $J_{w}\subseteq\operatorname{init}(I_{w})\subseteq\operatorname{init}(I(\mathsf{MSV}_{w}))$.
Theorem 4.3.
The set $\mathcal{T}=\{(I(\mathsf{MSV}_{w}),J_{w}):w\in\mathsf{Mat}_{m\times n}\}$ is a finite transition system.
We outline the proof of this theorem below. Before commencing this, we first explain some of the additional data that goes into this statement, in particular: (a) a finite indexing set for $\mathcal{T}$, (b) how to identify the ideals $I(\mathsf{MSV}_{w})$ that are maximal, (c) for non-maximal $I(\mathsf{MSV}_{w})$ how to construct the non-constant polynomial $f\notin I(\mathsf{MSV}_{w})$ and subset $\Phi\subseteq\mathcal{T}$ required in Definition 3.1.
4.2 Finite indexing set
Observe that $\mathsf{MSV}_{w}=\mathsf{MSV}_{gwh^{\top}}$, $I_{w}=I_{gwh^{\top}}$, and $J_{w}=J_{gwh^{\top}}$ for all $(g,h)\in\mathsf{B}_{m}\times\mathsf{B}_{n}$ and $w\in\mathsf{Mat}_{m\times n}$. Meanwhile, the $\mathsf{B}_{m}\times\mathsf{B}_{n}$-orbit of any element of $\mathsf{Mat}_{m\times n}$ contains a unique $m\times n$ partial permutation matrix [32, Ch. 15], so the size of $\mathcal{T}$ in Theorem 4.3 is at most the number of such matrices. The following notation is convenient to exploit these observations.
[…]
$S^{(m,n)}_{\infty}:=\{w\in S_{\infty}:\mathrm{Des}_{R}(w)\subseteq[m]\text{ and }\mathrm{Des}_{L}(w)\subseteq[n]\}$ (4.5)
is in bijection with the finite set of $m\times n$ partial permutation matrices, via the map that identifies $w\in S^{(m,n)}_{\infty}$ with the $m\times n$ matrix having $1$ in position $(i,w(i))$ for each $i\in[m]$ and 0 elsewhere. We freely make this identification in order to define $\mathsf{MSV}_{w}$, $I_{w}$, and $J_{w}$ for $w\in S^{(m,n)}_{\infty}$. Then
$\mathcal{T}=\left\{(I(\mathsf{MSV}_{w}),J_{w}):w\in\mathsf{Mat}_{m\times n}\right\}=\left\{(I(\mathsf{MSV}_{w}),J_{w}):w\in S^{(m,n)}_{\infty}\right\}$ (4.6)
and the set on the right is uniquely indexed, since if $v,w\in S^{(m,n)}_{\infty}$ then $I(\mathsf{MSV}_{v})\subseteq I(\mathsf{MSV}_{w})$ if and only if $v\leq w$ in the Bruhat order on $S_{\infty}$ [2, Thm. 2.1.5].
Example 4.4.
The ideal $I(\mathsf{MSV}_{w})$ is maximal if and only if $\mathsf{MSV}_{w}$ is a point, which occurs only when $\mathsf{MSV}_{w}=\{0\}$ as $\mathbb{K}^{\times}=\mathbb{K}\setminus\{0\}$ acts on $\mathsf{MSV}_{w}$ by multiplication. In this case the index is $w\in S_{\infty}^{(m,n)}$ is $0_{m\times n}:=(n+1)(n+2)\cdots(m+n)12\cdots n\in S_{\infty}$ and the corresponding ideals are irrelevant in the sense that $I(\mathsf{MSV}_{0_{m\times n}})=I_{0_{m\times n}}=J_{0_{m\times n}}=\operatorname{init}(I(\mathsf{MSV}_{0_{m\times n}}))=\langle x_{ij}:(i,j)\in[m]\times[n]\rangle.$
4.3 Proof of Theorem 4.3
[…]
where we compute $\operatorname{rank}w_{[i][j]}$ by identifying $w$ with the (infinite) permutation matrix having $1$ in each position $(i,w(i))$. The set $\mathsf{dom}(w)$ always coincides with the Young diagram
[…]
Example 4.5.
Consider $w=43152\in S_{5}$ written in one-line notation. Then $\mathrm{Des}_{R}(w)=\{1,2,4\}$ and $\mathrm{Des}_{L}(w)=\{2,3\}$ so $w\in S_{\infty}^{(m,n)}$ for any $m\geq 4$ and $n\geq 3$. The dominant component of $w$ is
$\mathsf{dom}(w)=\left\{\begin{smallmatrix}\square&\square&\square&1&\cdot\\ \square&\square&1&\cdot&\cdot\\ 1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1\\ \cdot&1&\cdot&\cdot&\cdot\end{smallmatrix}\right\}=\{(1,1),(1,2),(1,3),(2,1),(2,2)\}=\textsf{D}_{\lambda}$
for $\lambda=(3,2)$. The outer corners of $\mathsf{dom}(w)$ are $(1,4)$, $(2,3)$, and $(3,1)$.
Proposition 4.6.
Suppose $w\in S_{\infty}^{(m,n)}$ and $(i,j)\in[m]\times[n]$. Then $(i,j)\in\mathsf{dom}(w)$ if and only if $x_{ij}\in I(\mathsf{MSV}_{w})$ if and only if $x_{ij}\in J_{w}$.
Proof.
Assume $(i,j)\in\mathsf{dom}(w)$. Then $\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}=0$, so $M_{ij}=0$ for all $M\in\mathsf{MSV}_{w}$, and therefore $x_{ij}\in I(\mathsf{MSV}_{w})$. We also have $x_{ij}\in J_{w}$ as $x_{ij}=\operatorname{adiag}(\mathcal{X}_{RC})$ for $R=\{i\}$ and $C=\{j\}$.
Now suppose $(i,j)\notin\mathsf{dom}(w)$. Then $\operatorname{rank}w_{[k][l]}\geq\operatorname{rank}w_{[i][j]}\geq 1$ whenever $m\geq k\geq i$ and $n\geq l\geq j$. It follows that $\mathsf{MSV}_{w}$ contains the matrix $E_{ij}$ with $1$ in position $(i,j)$ and 0 in all other positions, so $x_{ij}\notin I(\mathsf{MSV}_{w})$ as $x_{ij}(E_{ij})=1\neq 0$. It also follows that $x_{ij}$ does not occur as any of the monomials $\operatorname{adiag}(\mathcal{X}_{RC})$ generating $J_{w}$ in (4.4), so $x_{ij}\notin J_{w}$ since $J_{w}$ is a monomial ideal. ∎
For positive integers $i<j$ let $t_{ij}=(i,j)\in S_{\infty}$ be the transposition swapping $i$ and $j$. We write $u\lessdot v$ for $u,v\in S_{\infty}$ if $v$ covers $u$ in the Bruhat order on $S_{\infty}$. Recall from [2, Lem. 2.1.4] that this holds precisely when $v=ut_{ij}$ for integers $i<j$ with $u(i)<u(j)$ such that no $i<e<j$ has $u(i)<u(e)<u(j)$.
Suppose $0_{m\times n}\neq w\in S_{\infty}^{(m,n)}$ so that $I(\mathsf{MSV}_{w})$ is not a maximal ideal. Then there exists an outer corner of $\mathsf{dom}(w)$ in $m]\times[n]$. Choose any such outer corner $(p,q)\in[m]\times[n]$ and set
$f=x_{pq}\in\mathbb{K}[\mathsf{Mat}_{m\times n}].$ (4.9)
Since $(p,q)\notin\mathsf{dom}(w)$, we have $f\notin I(\mathsf{MSV}_{w})$ by Proposition 4.6. Next, consider the set
[…]
It is known that $\varnothing\subsetneq\mathcal{C}(w)\subseteq S_{\infty}^{(m,n)}$ by [31, Lem. 5.8]. Finally define
$\Phi=\{(I(\mathsf{MSV}_{v}),J_{v}):v\in\mathcal{C}(w)\}.$ (4.11)
To make these definitions canonical, one can take $(p,q)$ to be the lexicographically minimal outer corner of $\mathsf{dom}(w)$ in $m]\times[n]$, although any outer corner works just as well.
Proof of Theorem 4.3.
We have already observed in Sections 4.1 and 4.2 that axioms (T1) and (T2) in Definition 3.1 hold.
For the other axioms, fix $w\in S_{\infty}^{(m,n)}$ with $I(\mathsf{MSV}_{w})$ not maximal. Choose an outer corner $(p,q)\in[m]\times[n]$ of $\mathsf{dom}(w)$, and define $f=x_{pq}\notin I(\mathsf{MSV}_{w})$ and $\Phi\subseteq\mathcal{T}$ as in (4.9) and (4.11). The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}_{w}),J_{w})$ are provided in [31, Lem. 5.11].
It remains to check axiom (T4). This holds by default as $(I(\mathsf{MSV}_{w}):\langle f\rangle)=I(\mathsf{MSV}_{w})$. There is a simple geometric reason for this equality: if $H\subset\mathsf{Mat}_{m\times n}$ is the hyperplane where $f=0$, then
$(I(\mathsf{MSV}_{w}):\langle f\rangle)=I(\overline{\mathsf{MSV}_{w}\setminus H}),$
and $\overline{\mathsf{MSV}_{w}\setminus H}=\mathsf{MSV}_{w}$ because $\mathsf{MSV}_{w}$ is an irreducible variety with $\mathsf{MSV}_{w}\not\subseteq H$ as $f\notin I(\mathsf{MSV}_{w})$.
One can also show $(I(\mathsf{MSV}_{w}):\langle f\rangle)=I(\mathsf{MSV}_{w})$ by an algebraic method based on Lemma 3.18. See the proof of Theorem 5.3 for a prototype of this argument. ∎
The preceding theorem lets us recover some results of Knutson and Miller [25].
Corollary 4.7 ([25]).
If $w\in\mathsf{Mat}_{m\times n}$ then $I_{w}=I(\mathsf{MSV}_{w})$ is prime and $J_{w}=\operatorname{init}(I_{w})$, and the set of minors generating $I_{w}$ in (4.3) is a Gröbner basis in the reverse lexicographic term order.
Proof.
Corollary 3.6 and Theorem 4.3 imply that the generating set in (4.3) is a Gröbner basis for $I(\mathsf{MSV}_{w})$. Thus $I_{w}=I(\mathsf{MSV}_{w})$ is prime as $\mathsf{MSV}_{w}$ is irreducible [32, Ch. 15], and $J_{w}=\operatorname{init}(I_{w})$. ∎
5 Skew-symmetric matrix Schubert varieties
One can construct a similar transition system for the skew-symmetric analogues of $\mathsf{MSV}_{w}$. We describe this here. This will generalize Example 3.7 and give another application of Theorem 3.3, recovering some results from [31].
Recall that $\mathbb{K}$ is an arbitrary field that is algebraically closed. We define a matrix $M$ to be skew-symmetric if $M^{\top}=-M$ and all diagonal entries of $M$ are zero; such matrices are also often called alternating. Let $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ be the affine variety of skew-symmetric $n\times n$ matrices over $\mathbb{K}$.
Definition 5.1.
For each skew-symmetric matrix $w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$, the corresponding skew-symmetric matrix Schubert cell and skew-symmetric matrix Schubert variety are the respective subsets
$\mathsf{MSC}^{\mathrm{ss}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (5.1)
$\mathsf{MSV}^{\mathrm{ss}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
where $M_{[i][j]}$ again stands for the upper left $i\times j$ submatrix of $M$.
Besides in [31], the varieties $\mathsf{MSV}^{\mathrm{ss}}_{w}$ were previously studied in [18] and [28], where their cohomology and $K$-theory classes were respectively computed (for relevant formulas, see also [29, 30]). The varieties $\mathsf{MSV}^{\mathrm{ss}}_{w}$ include a number of other families as special cases, like all rank $r$ skew-symmetric $n\times n$ matrices. Questions about ideals and Gröbner bases for some of these families were considered in various forms in earlier literature; we note in particular [8, 9, 21, 23, 33].
We identify the coordinate ring of $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ with $\mathbb{K}[\mathsf{Mat}^{\mathrm{ss}}_{n\times n}]=\mathbb{K}[u_{ij}:1\leq j<i\leq n]$ where $u_{ij}$ denotes the linear function on matrices with $u_{ij}(M)=M_{ij}\in\mathbb{K}$. We order the monomials in this ring using the reverse lexicographic term order explained in Remark 2.3.
5.1 Another reverse lexicographic transition system
Fix a matrix $w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$. It is known from [4] that the variety $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is irreducible and equal to the Zariski closure of $\mathsf{MSC}^{\mathrm{ss}}_{w}$. The group $\mathsf{B}_{n}$ of invertible invertible lower-triangular $n\times n$ matrices over $\mathbb{K}$ acts algebraically on $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ by $g\cdot M=gMg^{\top}$, and the skew-symmetric matrix Schubert cell $\mathsf{MSC}^{\mathrm{ss}}_{w}$ is the orbit of $w$ under this action. Thus, in the notation of Section 3.3, we can express
$I(\mathsf{MSV}^{\mathrm{ss}}_{w})=\mathrm{cl}_{\mathsf{B}_{n}}(\mathcal{M}^{\mathrm{ss}}_{w})\quad\text{for the maximal ideal }\mathcal{M}^{\mathrm{ss}}_{w}:=\langle u_{ij}-w_{ij}:1\leq j<i\leq n\rangle.$ (5.2)
Define $\mathcal{U}^{\mathrm{ss}}$ to be the $n\times n$ skew-symmetric matrix with entries $\mathcal{U}^{\mathrm{ss}}_{ij}=u_{ij}=-\mathcal{U}^{\mathrm{ss}}_{ji}$ for $i>j$ and with $\mathcal{U}^{\mathrm{ss}}_{ii}=0$ for all $i$. Then it is true, just as for ordinary matrix Schubert varieties, that
$I(\mathsf{MSV}^{\mathrm{ss}}_{w})\supseteq\left\langle\det(\mathcal{U}^{\mathrm{ss}}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle.$ (5.3)
This implies that, relative to the reverse lexicographic term order, one has
$\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))\supseteq\left\langle\operatorname{adiag}(\mathcal{U}^{\mathrm{ss}}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle.$ (5.4)
Unlike the story for ordinary matrix Schubert varieties, both of these containments can be strict [31, Ex. 3.12]. Therefore, to get a skew-symmetric version of Theorem 4.3, we cannot use the right hand side of (5.4) as a skew-symmetric analogue of $J_{w}$.
Instead, we consider a more elaborate construction from [31]. Suppose $A=\{a_{0}<a_{1}<\dots<a_{r}\}$ and $B=\{b_{0}<b_{1}<\dots<b_{r}\}$ are two sets of $r+1$ positive integers. Define
$A\odot B=\{(a_{0},b_{r}),(a_{1},b_{r-1}),\dots,(a_{r},b_{0})\}=\{(a_{i},b_{j}):i+j=r\}.$ (5.5)
Then let $A\boxplus B:=\{(i,j)\in A\odot B:i\geq j\}\cup\{(i,j)\in B\odot A:i\geq j\}$ and set
$u^{\mathrm{ss}}_{AB}:=\begin{cases}\prod_{(i,j)\in A\boxplus B}u_{ij}&\text{if every $(i,j)\in A\boxplus B$ has $i\neq j$}\\ 0&\text{otherwise}.\end{cases}$ (5.6)
For example, if $a\neq b$ then $u^{\mathrm{ss}}_{\{a\}\{b\}}=u_{\max\{a,b\}\min\{a,b\}}$ while if $A=\{1,3,4\}$ and $B=\{2,5,6\}$ then $A\boxplus B=\{(4,2),(5,3),(6,1)\}$ and $u^{\mathrm{ss}}_{AB}=u_{42}u_{53}u_{61}.$ When nonzero, the monomial $u^{\mathrm{ss}}_{AB}$ is always square-free. We record a lemma for later use:
Lemma 5.2.
If $u^{\mathrm{ss}}_{AB}=u_{ij}$ for some $n\geq i>j\geq 1$ then $A=B=\{i,j\}$ or $\{A,B\}=\{\{i\},\{j\}\}$.
Proof.
This can be shown in a self-contained way, but the elementary argument is fairly tedious. Alternatively, a special case of [31, Lem 3.25] asserts that $u^{\mathrm{ss}}_{AB}$ divides $u_{ij}$ if and only if $\operatorname{adiag}(\mathcal{X}_{AB})$ divides $x_{ij}x_{ji}$, which occurs either when $A=B=\{i,j\}$ or $\{A,B\}=\{\{i\},\{j\}\}$. ∎
Now we introduce the monomial ideal
$J^{\mathrm{ss}}_{w}=\left\langle u^{\mathrm{ss}}_{AB}:i,j\in[n],\ i\geq j,\ A\subseteq[i],\ B\subseteq[j],\ |A|=|B|=1+\operatorname{rank}w_{[i][j]}\right\rangle.$ (5.7)
Unlike for the ideals in Section 4, it is not obvious that $J^{\mathrm{ss}}_{w}\subseteq\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))$. Nevertheless:
Theorem 5.3.
The set $\mathcal{T}^{\mathrm{ss}}=\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}\}$ is a finite transition system.
We explain the proof of this theorem below, following a strategy similar to the one in Section 4.
5.2 Finite indexing set
If $w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ then $\mathsf{MSV}^{\mathrm{ss}}_{w}=\mathsf{MSV}^{\mathrm{ss}}_{gwg^{\top}}$ and $J^{\mathrm{ss}}_{w}=J^{\mathrm{ss}}_{gwg^{\top}}$ for all $g\in\mathsf{B}_{n}$, so the set $\mathcal{T}^{\mathrm{ss}}$ has cardinality at most the number $\mathsf{B}_{n}$-orbits in $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$. We recall a finite indexing set for these orbits.
[…]
For each $w\in\mathcal{I}_{\textsf{FPF}}$ define $\mathrm{ss}_{n}(w)\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ to be the matrix whose entry in position $(i,j)$ is $1$ if $i<j=w(i)$, 1-1 if $i>j=w(i)$, or else zero. Then each $\mathsf{B}_{n}$-orbit in $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ contains an element of the form $\mathrm{ss}_{n}(w)$ for a unique $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ by results in [4]; it is not hard to show that this holds even when $\mathbb{K}$ is not algebraically closed. Given $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ we are therefore motivated to define
$\mathsf{MSC}^{\mathrm{ss}}_{w}:=\mathsf{MSC}^{\mathrm{ss}}_{\mathrm{ss}_{n}(w)},\qquad\mathsf{MSV}^{\mathrm{ss}}_{w}:=\mathsf{MSV}^{\mathrm{ss}}_{\mathrm{ss}_{n}(w)},\qquad\text{and}\qquad J^{\mathrm{ss}}_{w}:=J^{\mathrm{ss}}_{\mathrm{ss}_{n}(w)}.$ (5.9)
Observe that $\mathsf{MSC}^{\mathrm{ss}}_{w}=\mathring{X}_{w}\cap\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ and $\mathsf{MSV}^{\mathrm{ss}}_{w}=X_{w}\cap\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$. Finally, we can write
$\mathcal{T}^{\mathrm{ss}}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathcal{I}_{\textsf{FPF}}^{(n)}\right\}$
and the second set is uniquely indexed, since if $v,w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{ss}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $v\leq w$ in a certain Bruhat order on $\mathcal{I}_{\textsf{FPF}}$ [31, Prop. 3.11].
Example 5.4.
As in the classical case, the ideal $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{ss}}_{w}=\{0\}$. The index $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ for this case is
$0^{\mathrm{ss}}_{n\times n}:=(1,n+1)(2,n+2)\cdots(n,2n)(2n+1,2n+2)(2n+3,2n+4)\cdots\in\mathcal{I}_{\textsf{FPF}}^{(n)}$
and the ideals $I(\mathsf{MSV}^{\mathrm{ss}}_{0^{\mathrm{ss}}_{n\times n}})=J^{\mathrm{ss}}_{0^{\mathrm{ss}}_{n\times n}}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{0^{\mathrm{ss}}_{n\times n}}))=\langle u_{ij}:i,j\in[n],\ i>j\rangle$ are all irrelevant.
5.3 Proof of Theorem 5.3
Recall the definition of $\mathsf{dom}(w)$ and its set of outer corners from Section 4.3.
Proposition 5.5.
Suppose $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ and $(i,j)\in[n]\times[n]$ has $i>j$. Then $(i,j)\in\mathsf{dom}(w)$ if and only if $u_{ij}\in I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $u_{ij}\in J^{\mathrm{ss}}_{w}$.
Proof.
Our argument is only slightly more complicated than the proof of Proposition 4.6. First, assume $(i,j)\in\mathsf{dom}(w)$. Then $\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}=0$ and so $M_{ij}=0$ for all $M\in\mathsf{MSV}^{\mathrm{ss}}_{w}$, which means that $u_{ij}\in I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. We also have $u_{ij}\in J^{\mathrm{ss}}_{w}$ since $u_{ij}=u^{\mathrm{ss}}_{AB}$ for $A=\{i\}$ and $B=\{j\}$.
Suppose conversely that $(i,j)\notin\mathsf{dom}(w)$. As the partial permutation matrix of $w$ is symmetric with zeros on the diagonal, it follows that $\operatorname{rank}w_{[k][l]}=\operatorname{rank}w_{[l][k]}\geq\operatorname{rank}w_{[i][j]}\geq 1$ whenever $n\geq k\geq i$ and $n\geq l\geq j$, and also that $\operatorname{rank}w_{[k][l]}\geq\operatorname{rank}w_{[i][i]}\geq 2$ whenever $\min\{k,l\}\geq i$. One concludes that $\mathsf{MSV}^{\mathrm{ss}}_{w}=X_{w}\cap\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ contains the matrix $E_{ij}-E_{ji}$ so $u_{ij}\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$.
To show that $u_{ij}\notin J^{\mathrm{ss}}_{w}$, we just need to check that $u_{ij}$ does not equal any of the terms $u^{\mathrm{ss}}_{AB}$ generating $J^{\mathrm{ss}}_{w}$ in (5.7). This follows from Lemma 5.2, as if $A\subseteq[k]$ and $B\subseteq[l]$ have $A=B=\{i,j\}$, then $\min\{k,l\}\geq i$ so $\operatorname{rank}w_{[k][l]}\geq 2$, while if $\{A,B\}=\{\{i\},\{j\}\}$ then similarly $\operatorname{rank}w_{[k][l]}\geq 1$. ∎
Following [17, §4.1], we define the fpf-involution length of $w\in\mathcal{I}_{\textsf{FPF}}$ to be
[…]
Write $u\lessdot_{\textsf{FPF}}v$ for $u,v\in\mathcal{I}_{\textsf{FPF}}$ if $v=t_{ij}\cdot u\cdot t_{ij}$ for any positive integers $i<j$ with $\ell_{\textsf{FPF}}(v)=\ell_{\textsf{FPF}}(u)+1$. For a description of $\lessdot_{\textsf{FPF}}$ in terms of the cycles of $u$ and $v$, see [17, Prop. 4.9].
Suppose $0^{\mathrm{ss}}_{n\times n}\neq w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ so that $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is not a maximal ideal and $\mathsf{MSV}^{\mathrm{ss}}_{w}\neq\{0\}$. Then $\mathsf{dom}(w)$, which is invariant under transpose, cannot contain every position in $n]\times[n]$, so there must exist an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p\geq q\geq 1$. As explained in [31, §4.1], all outer corners $(p,q)$ of $\mathsf{dom}(w)$ have $w(p)=q$, so as $w$ has no fixed points we can assume that $p>q$.
In summary, we can choose an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p>q\geq 1$. Make such a choice (or to be canonical, let $(p,q)$ be the lexicographically minimal choice) and then set
$f=u_{pq}\in\mathbb{K}[\mathsf{Mat}^{\mathrm{ss}}_{m\times n}].$ (5.11)
Since $(p,q)\notin\mathsf{dom}(w)$, we have $f\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ by Proposition 5.5. Next, define the set
[…]
It is known that $\varnothing\subsetneq\mathcal{C}^{\mathrm{ss}}(w)\subseteq\mathcal{I}_{\textsf{FPF}}^{(n)}$ by [31, Lem. 4.8]. Finally, let
$\Phi=\{(I(\mathsf{MSV}^{\mathrm{ss}}_{v}),J^{\mathrm{ss}}_{v}):v\in\mathcal{C}^{\mathrm{ss}}(w)\}.$ (5.13)
Proof of Theorem 5.3.
The claim that $J^{\mathrm{ss}}_{w}\subseteq\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))$ holds by [31, Lem 3.37 and Thm. 3.17]. The other parts of axioms (T1) and (T2) in Definition 3.1 were checked in Sections 5.2 and 5.3.
Fix $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ with $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ not maximal, and then choose an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p>q\geq 1$. The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w})$ are supplied in [31, Lem. 4.2] if we define $f=u_{pq}\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ and $\Phi$ as in (5.11) and (5.13).
It will turn out that axiom (T4) holds vacuously as $(I(\mathsf{MSV}^{\mathrm{ss}}_{w}):\langle f\rangle)=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. To show this, we examine the group action of $G=\mathsf{B}_{n}$ on $A=\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$. If we identify
$\mathbb{K}[\mathsf{B}_{n}]=\mathbb{K}\left[b_{ij},b_{ii}^{-1}:1\leq j\leq i\leq n\right]$
where $b_{ij}(g)=g_{ij}$, then the map $\alpha^{\ast}$ from (3.6) has the formula
$\alpha^{\ast}(u_{ij})=\sum_{\begin{subarray}{c}p,q\in[n]\\ 1\leq q\leq j<p\leq i\end{subarray}}b_{ip}b_{jq}\otimes u_{pq}+\sum_{\begin{subarray}{c}p,q\in[n]\\ 1\leq q<p\leq j<i\end{subarray}}(b_{ip}b_{jq}-b_{iq}b_{jp})\otimes u_{pq}\quad\text{for $1\leq j<i\leq n$.}$ (5.14)
The hypothesis of Lemma 3.18 holds for the partial order $\preceq$ on $\mathsf{Index}:=\{(i,j)\in[n]\times[n]:i>j\}$ with $(i,j)\preceq(k,l)$ if both $i\leq k$ and $j\leq l$. Since $(p,q)$ is an outer corner of $\mathsf{dom}(w)$, we have $(i,j)\in\mathsf{dom}(w)$ for all $(i,j)\in\mathsf{Index}$ with $(i,j)\prec(p,q)$, and so $u_{ij}\in I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ for all $(i,j)\in\mathsf{Index}$ with $(i,j)\prec(p,q)$ by Proposition 5.5. Thus, referring to (5.2), we conclude by Lemma 3.18(a) that $u_{pq}\notin\mathcal{M}^{\mathrm{ss}}_{w}$ since $u_{pq}\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. Finally, from Lemmas 3.18(b) and 3.19, we get
$(I(\mathsf{MSV}^{\mathrm{ss}}_{w}):\langle f\rangle)=(\mathrm{cl}_{\mathsf{B}_{n}}(\mathcal{M}^{\mathrm{ss}}_{w}):\langle u_{pq}\rangle)=\mathrm{cl}_{\mathsf{B}_{n}}((\mathcal{M}^{\mathrm{ss}}_{w}:\langle u_{pq}\rangle))=\mathrm{cl}_{\mathsf{B}_{n}}(\mathcal{M}^{\mathrm{ss}}_{w})=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$
as promised, and so axiom (T4) is automatically satisfied. ∎
Theorem 5.3 immediately recovers the following property from [31, Thm. 4.3].
Corollary 5.6 ([31]).
If $w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ then $J^{\mathrm{ss}}_{w}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))$.
So far we have avoided defining a skew-symmetric analogue of the ideal $I_{w}$ from (4.3). Such an ideal $I^{\mathrm{ss}}_{w}\subseteq\mathbb{K}[\mathsf{Mat}^{\mathrm{ss}}_{n\times n}]$ is introduced for each $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ in [31, Def. 3.13]. We omit an explicit description here, except to state that this ideal $I^{\mathrm{ss}}_{w}$ can be generated a set of Pfaffians of certain submatrices of the skew-symmetric matrix of variables $\mathcal{U}^{\mathrm{ss}}=\begin{bmatrix}0&-u_{ij}\\ u_{ij}&0\end{bmatrix}$.
The simplest Pfaffian generating set for $I^{\mathrm{ss}}_{w}$ given in [31, Def. 3.13] is not a Gröbner basis [31, Ex. 3.28]. However, there is a more complicated set of Pfaffians $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}$ described in [31, Thm. 4.6] for which it turns out that $J^{\mathrm{ss}}_{w}\subseteq\langle\operatorname{init}(g):g\in G_{w}^{\mathrm{ss}}\rangle$ [31, Lems. 3.35 and 3.36].
It can be shown that $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is the zero locus of $I^{\mathrm{ss}}_{w}$ and so $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ [31, Thm. 3.17]. Therefore, Corollary 3.6 implies that $G_{w}^{\mathrm{ss}}$ is a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$, so $I^{\mathrm{ss}}_{w}=\langle G_{w}^{\mathrm{ss}}\rangle=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is prime (as $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is irreducible) and $J^{\mathrm{ss}}_{w}=\operatorname{init}(I^{\mathrm{ss}}_{w})$. These facts recover [31, Thms. 4.5 and 4.6].
6 Symmetric matrix Schubert varieties
[…]
Definition 6.1.
For each symmetric matrix $w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$, the corresponding symmetric matrix Schubert cell and symmetric matrix Schubert variety are the respective subsets
$\mathsf{MSC}^{\mathrm{sym}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (6.1)
$\mathsf{MSV}^{\mathrm{sym}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
where as usual $M_{[i][j]}$ stands for the upper left $i\times j$ submatrix of $M$.
As with $\mathsf{MSV}^{\mathrm{ss}}_{w}$, the varieties $\mathsf{MSV}^{\mathrm{sym}}_{w}$ include many familiar classes of symmetric matrices, whose ideals have been previously studied (see, e.g., [15, 27]). Questions about the Gröbner geometry of symmetric matrix Schubert varieties are less well-understood. We know of only a few relevant references [6, 7, 10, 16], which supply answers in special cases.
Remark.
Ideals and Gröbner bases for a different collection of “symmetric matrix Schubert varieties” are considered in [11, 12]. The varieties in these references are defined by imposing northeast rank conditions, while $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is defined via northwest rank conditions. The two families are not generally related in any simple way.
We identify the coordinate ring of $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ with $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]=\mathbb{K}[u_{ij}:1\leq j\leq i\leq n]$ where $u_{ij}$ represents the linear function on matrices with $u_{ij}(M)=M_{ij}\in\mathbb{K}$. We continue to order the monomials in this ring using the reverse lexicographic term order explained in Remark 2.3. (Note that this is not the term order considered in [6].)
It is an open problem to determine the initial ideals and associated Gröbner bases for $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. Computations support a plausible conjecture. Below, we discuss this conjecture along with a speculative proof strategy using transition systems. Turning this approach into a detailed proof will require new ideas to overcome obstacles that did not arise in Sections 4 or 5.
6.1 An incomplete transition system
[…]
Remark 6.2.
The lower-triangular Borel subgroup $\mathsf{B}_{n}$ acts algebraically on $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ by the same formula as in the skew-symmetric case: $g\cdot M=gMg^{\top}$. When $\operatorname{\mathsf{char}}(\mathbb{K})\neq 2$, the matrix Schubert cell $\mathsf{MSC}^{\mathrm{sym}}_{w}$ is the orbit of $w$ under this action, and so in the notation of Section 3.3, we have
$I(\mathsf{MSV}^{\mathrm{sym}}_{w})=\mathrm{cl}_{\mathsf{B}_{n}}(\mathcal{M}^{\mathrm{sym}}_{w})\quad\text{for the maximal ideal }\mathcal{M}^{\mathrm{sym}}_{w}:=\langle u_{ij}-w_{ij}:1\leq j\leq i\leq n\rangle.$ (6.2)
However, if $\operatorname{\mathsf{char}}(\mathbb{K})=2$ then $\mathsf{MSC}^{\mathrm{sym}}_{w}$ may be a union of multiple of $\mathsf{B}_{n}$-orbits.
Define $\mathcal{U}^{\mathrm{sym}}$ to be the $n\times n$ symmetric matrix with entries $\mathcal{U}^{\mathrm{sym}}_{ij}=u_{ij}=\mathcal{U}^{\mathrm{sym}}_{ji}$ for all $i\geq j$. Similar to our two previous cases, if we consider the ideal generated by minors
$I^{\mathrm{sym}}_{w}:=\left\langle\det(\mathcal{U}^{\mathrm{sym}}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle$ (6.3)
and also define
$J^{\mathrm{sym}}_{w}:=\left\langle\operatorname{adiag}(\mathcal{U}^{\mathrm{sym}}_{RC}):i\in[m],\ j\in[n],\ R\subseteq[i],\ C\subseteq[j],\ |R|=|C|=1+\operatorname{rank}w_{[i][j]}\right\rangle$ (6.4)
then it automatically holds that
$I^{\mathrm{sym}}_{w}\subseteq I(\mathsf{MSV}^{\mathrm{sym}}_{w})\quad\text{and}\quad J^{\mathrm{sym}}_{w}\subseteq\operatorname{init}(I^{\mathrm{sym}}_{w})\subseteq\operatorname{init}(I(\mathsf{MSV}^{\mathrm{sym}}_{w}))$ (6.5)
relative to the reverse lexicographic term order.
We expect that the two containments in (6.5) are actually both equalities (see Conjectures 6.5 and 6.6), just as in the ordinary matrix Schubert case. However, this cannot be shown by proving the most obvious symmetric reformulation of Theorem 4.3, since the set
$\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\}$ (6.6)
is generally too small to be a transition system (or even a partial transition system), as we have already seen in Example 3.9. Nevertheless, computations support the following conjecture:
Conjecture 6.3.
There exists a transition system containing $\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\}$.
As observed in Example 3.9, the transition systems realizing Conjecture 6.3 will have to include non-radical ideals of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$. We expect that these systems will still be finite, however.
One can see that the set $\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\}$ is itself finite by noting that the set is unchanged when we restrict $w$ to range over all symmetric $n\times n$ partial permutation matrices, since these matrices give rise to all possible rank tables for elements of $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. The latter claim can be deduced (when $\mathbb{K}$ is an arbitrary field) from [28, Lem. 3.20], which shows that the rank table of any matrix coincides with the rank table of some partial permutation matrix, and the latter must be symmetric for its rank table to be symmetric.
We will index the symmetric $n\times n$ partial permutation matrices by setting
$\mathcal{I}_{\infty}:=\{w\in S_{\infty}:w=w^{-1}\}\quad\text{and}\quad\mathcal{I}_{\infty}^{(n)}:=\{w\in\mathcal{I}_{\infty}:\mathrm{Des}_{R}(w)\subseteq[n]\}=\mathcal{I}_{\infty}\cap S_{\infty}^{(n,n)}.$ (6.7)
Passing to the $n\times n$ partial permutation matrix gives a bijection from $\mathcal{I}_{\infty}^{(n)}$ to symmetric $n\times n$ partial permutation matrices. This lets us define $\mathsf{MSV}^{\mathrm{sym}}_{w}$, $I^{\mathrm{sym}}_{w}$, and $J^{\mathrm{sym}}_{w}$ for $w\in\mathcal{I}_{\infty}^{(n)}$. Then
$\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(n)}\right\},$ (6.8)
and the second set is uniquely indexed since if $v,w\in\mathcal{I}_{\infty}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{sym}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ if and only if $v\leq w$ in the Bruhat order of $S_{\infty}$ by [1, Lem. 3.5].
Example 6.4.
The ideal $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{sym}}_{w}=\{0\}$. The index $w\in\mathcal{I}_{\infty}^{(n)}$ for this case is $0_{n\times n}=(1,n+1)(2,n+2)\cdots(n,2n)$, and we have $I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}})=I^{\mathrm{sym}}_{0_{n\times n}}=J^{\mathrm{sym}}_{0_{n\times n}}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}}))=\langle u_{ij}:i,j\in[n],\ i\geq j\rangle.$
We mention some interesting consequences of Conjecture 6.3. First, via Theorem 3.3, the conjecture would immediately imply that $J^{\mathrm{sym}}_{w}$ is the initial ideal of $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. Then it would follow by Corollary 3.6 that the minors listed in (6.3) form a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. As these minors already generate $I^{\mathrm{sym}}_{w}$, this would prove the following:
Conjecture 6.5.
If $w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ then $I^{\mathrm{sym}}_{w}=I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ and $J^{\mathrm{sym}}_{w}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{sym}}_{w}))$.
The next conjecture would also be a consequence:
Conjecture 6.6.
If $w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ is any symmetric matrix then $\operatorname{init}(I^{\mathrm{sym}}_{w})=J^{\mathrm{sym}}_{w}$ and the set of minors generating $I^{\mathrm{sym}}_{w}$ in (6.3) is a Gröbner basis in the reverse lexicographic term order.
Finally, as $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is irreducible, Conjecture 6.5 would imply this last property:
Conjecture 6.7.
If $w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ is any symmetric matrix then $I^{\mathrm{sym}}_{w}$ is a prime ideal of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$.
[…]
One application of the Gröbner basis computations for matrix Schubert varieties and their skew-symmetric counterparts is to determine primary decompositions of the initial ideals $J_{w}$ and $J^{\mathrm{ss}}_{w}$ for $I_{w}=I(\mathsf{MSV}_{w})$ and $I^{\mathrm{ss}}_{w}=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. These decompositions are intersections of monomial ideals indexed by certain pipe dreams of $w$; see [25, Thm. B] and [28, Thm. 4.15].
[…]
Computations for $n\leq 10$ support the following conjecture. Recall that $\mathcal{I}_{n}\subsetneq\mathcal{I}_{\infty}^{(n)}$ is the set of involutions in $S_{n}$, corresponding to the symmetric $n\times n$ partial permutation matrices of full rank.
Conjecture 6.8.
If $w\in\mathcal{I}_{n}$ then $J^{\mathrm{sym}}_{w}=\bigcap_{D}J^{\mathrm{sym}}_{D}$ where $J^{\mathrm{sym}}_{D}:=\left\langle u_{ij}^{m_{D}(i,j)}:(i,j)\in D\right\rangle$ and where $D$ runs over all involution pipe dreams for all $z\in\mathcal{I}_{n}$ with $z\geq w$ in Bruhat order.
Replacing $\mathcal{I}_{n}$ by $\mathcal{I}_{\infty}^{(n)}$ in this conjecture would give the most natural symmetric analogue of [31, Thm. 4.15]. However, this broader claim is false, and to decompose $J^{\mathrm{sym}}_{w}$ when $w$ corresponds to a symmetric $n\times n$ partial permutation matrix that is not invertible, a more involved statement is required.
We also mention that when $w\in\mathcal{I}_{n}$, the decomposition given in Conjecture 6.8 is very redundant. It suffices to let $D$ run over the minimal elements of the poset on involution pipe dreams for $z\geq w$ defined by $D\leq E$ if $J^{\mathrm{sym}}_{D}\subseteq J^{\mathrm{sym}}_{E}$. If the exponents $m_{D}(i,j)$ in the definition of $J^{\mathrm{sym}}_{D}$ were replaced by $1$, then these minimal elements would be just the involution pipe dreams for $w$. This does seem to hold if $w$ is non-crossing in the sense of being $3412$-avoiding, but is not true in general.
6.2 Transition forests
One can attempt to find a transition system containing a given family of ideals by recursively determining primary decompositions in the following way. This method is amenable to computer calculations using algebra systems like Macaulay2.
For a non-maximal ideal $I\subset\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$, let $u_{I}$ be the lexicographically smallest variable $u_{ij}\notin I$ and define $\mathscr{P}_{I}$ to be any set of primary ideals such that $I+\langle u_{ij}\rangle=\bigcap_{K\in\mathscr{P}_{I}}K$ is a primary decomposition. When the ideal $I$ is maximal, we leave $u_{I}$ undefined and set $\mathscr{P}_{I}=\varnothing$.
Now, from a given set of ideals $\mathscr{I}$, we form $\mathcal{T}_{\mathscr{I}}:=\mathcal{T}_{\mathscr{I}}^{0}\cup\mathcal{T}_{\mathscr{I}}^{1}\cup\mathcal{T}_{\mathscr{I}}^{2}\cup\cdots$ where
$\mathcal{T}_{\mathscr{I}}^{0}:=\{(I,\operatorname{init}(I)):I\in\mathscr{I}\}\quad\text{and}\quad\mathcal{T}_{\mathscr{I}}^{i+1}:=\left\{(K,\operatorname{init}(K)):K\in\bigcup_{(I,J)\in\mathcal{T}_{\mathscr{I}}^{i}}\mathscr{P}_{I}\right\}.$
This definition yields an increasing chain of subsets $\mathcal{T}_{\mathscr{I}}^{0}\subseteq\mathcal{T}_{\mathscr{I}}^{1}\subseteq\mathcal{T}_{\mathscr{I}}^{2}\subseteq\cdots$ which terminates in a finite number of steps since the ambient coordinate ring is Noetherian.
It is not guaranteed that this process will yield a family satisfying every part of Definition 3.1: problems can occur with conditions (T3) and (T4). However, the construction of $\mathcal{T}_{\mathscr{I}}$ sometimes does produce a transition system.
We visualize $\mathcal{T}_{\mathscr{I}}$ by drawing a forest graph whose vertices are the ideals $I$ with $(I,J)\in\mathcal{T}_{\mathscr{I}}$. In this transition forest, the children of a vertex labeled $I$ are the elements of $\mathscr{P}_{I}$. Although we have formulated these definitions specifically for the case $A=\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$, the relevant ideas easily extend to the case when $A$ is any affine space.
Example 6.9.
We revisit the ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(2)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{2\times 2}]$ from Example 3.9. If we define $I_{1},\dots,I_{6}$ as in that example, then $\mathscr{I}=\{I_{i}:i\in[5]\}$ and the transition forest of $\mathcal{T}_{\mathscr{I}}$ is
[…]
Here and in later pictures, we place boxes around the ideals $\boxed{I}$ with $I\notin\mathscr{I}$. As we know from Example 3.9, the set $\mathcal{T}_{\mathscr{I}}$ is a transition system and therefore verifies Conjecture 6.3 when $n=2$.
Recall that the associated primes of an ideal $I$ are the prime ideals $\left\{\sqrt{K_{1}},\sqrt{K_{2}},\ldots,\sqrt{K_{m}}\right\}$ where $I=K_{1}\cap K_{2}\cap\cdots\cap K_{m}$ is a primary decomposition. An embedded prime is an associated prime which is not minimal with respect to set-theoretic inclusion. If there are no embedded primes, then there is a unique primary decomposition $I=K_{1}\cap K_{2}\cap\cdots\cap K_{m}$ in which no $K_{i}$ is redundant and with all $\sqrt{K_{1}},\sqrt{K_{2}},\ldots,\sqrt{K_{m}}$ distinct.
If no embedded primes occur for the ideals $I+\langle u_{I}\rangle$ with $(I,J)\in\mathcal{T}_{\mathscr{I}}$, then the set $\mathcal{T}_{\mathscr{I}}$ is uniquely determined as long as we insist that each $\mathscr{P}_{I}$ be a unique irredundant primary decomposition. In general, however, there can be multiple ways of constructing $\mathcal{T}_{\mathscr{I}}$ from a given set of ideals $\mathscr{I}$. This phenomenon did not arise in the previous example but does in the following one.
Example 6.10.
Consider the ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(3)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$. In this case $\mathscr{I}$ has 14 elements $I_{1},I_{2},\dots,I_{14}$, which are indicated in Figure 1. We have used Macaulay2 to find one valid construction of $\mathcal{T}_{\mathscr{I}}$. This requires adding 7 more ideals $I_{15},I_{16},\dots,I_{21}$, also shown in Figure 1. Figure 2 displays the associated transition forest.
Embedded primes occur in this example, so $\mathcal{T}_{\mathscr{I}}$ is not unique. Specifically, for
$I_{19}=\left\langle{u}_{11},\ {u}_{32}^{2}-{u}_{22}{u}_{33},\ {u}_{31}{u}_{32}-{u}_{21}{u}_{33},\ {u}_{31}^{2},\ {u}_{22}{u}_{31}-{u}_{21}{u}_{32},\ {u}_{21}{u}_{31},\ {u}_{21}^{2}\right\rangle$
we have a primary decomposition $I_{19}+\langle u_{21}\rangle=I_{11}\cap I_{21}$ where
$I_{11}=\left\langle{u}_{31},\ {u}_{21},\ {u}_{11},\ {u}_{32}^{2}-{u}_{22}{u}_{33}\right\rangle\quad\text{and}\quad I_{21}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11},\ {u}_{32}^{2},\ {u}_{31}{u}_{32},\ {u}_{31}^{2}\right\rangle.$
However, it holds that $\sqrt{I_{11}}=I_{11}=\langle u_{31},u_{21},u_{11},u_{22}u_{33}-u_{32}^{2}\rangle\subsetneq\sqrt{I_{21}}=\langle u_{22},u_{21},u_{11},u_{31},u_{32}\rangle.$ Regardless, we have checked that the construction of $\mathcal{T}_{\mathscr{I}}$ corresponding to what is shown in Figure 2 is a transition system. This verifies Conjecture 6.3 when $n=3$.
$w\in\mathcal{I}_{\infty}^{(3)}$ $I(\mathsf{MSV}^{\mathrm{sym}}_{w})\subset\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$
$\left[\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{smallmatrix}\right]$ $I_{1}=0$
$\left[\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{smallmatrix}\right]$ $I_{2}=\left\langle{u}_{22}{u}_{31}^{2}-2{u}_{21}{u}_{31}{u}_{32}+{u}_{11}{u}_{32}^{2}+{u}_{21}^{2}{u}_{33}-{u}_{11}{u}_{22}{u}_{33}\right\rangle$
$\left[\begin{smallmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{smallmatrix}\right]$ $I_{3}=\left\langle{u}_{21}^{2}-{u}_{11}{u}_{22}\right\rangle$
$\left[\begin{smallmatrix}1&0&0\\ 0&0&0\\ 0&0&1\end{smallmatrix}\right]$ $I_{4}=\left\langle{u}_{22}{u}_{31}-{u}_{21}{u}_{32},\ {u}_{21}{u}_{31}-{u}_{11}{u}_{32},\ {u}_{21}^{2}-{u}_{11}{u}_{22}\right\rangle$
$\left[\begin{smallmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{smallmatrix}\right]$ $I_{5}=\left\langle\begin{array}[]{lll}{u}_{32}^{2}-{u}_{22}{u}_{33},&{u}_{31}{u}_{32}-{u}_{21}{u}_{33},&{u}_{31}^{2}-{u}_{11}{u}_{33},\\ {u}_{22}{u}_{31}-{u}_{21}{u}_{32},&{u}_{21}{u}_{31}-{u}_{11}{u}_{32},&{u}_{21}^{2}-{u}_{11}{u}_{22}\end{array}\right\rangle$
$\left[\begin{smallmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{smallmatrix}\right]$ $I_{6}=\left\langle{u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{smallmatrix}\right]$ $I_{7}=\left\langle{u}_{11},\ {u}_{22}{u}_{31}^{2}-2{u}_{21}{u}_{31}{u}_{32}+{u}_{21}^{2}{u}_{33}\right\rangle$
$\left[\begin{smallmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{smallmatrix}\right]$ $I_{8}=\left\langle{u}_{21},\ {u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{smallmatrix}\right]$ $I_{9}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&0&0\\ 0&1&0\\ 0&0&1\end{smallmatrix}\right]$ $I_{10}=\left\langle{u}_{31},\ {u}_{21},\ {u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{smallmatrix}\right]$ $I_{11}=\left\langle{u}_{31},\ {u}_{21},\ {u}_{11},\ {u}_{32}^{2}-{u}_{22}{u}_{33}\right\rangle$
$\left[\begin{smallmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{smallmatrix}\right]$ $I_{12}=\left\langle{u}_{31},\ {u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{smallmatrix}\right]$ $I_{13}=\left\langle{u}_{32},\ {u}_{31},\ {u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$
$\left[\begin{smallmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{smallmatrix}\right]$ $I_{14}=\left\langle{u}_{33},\ {u}_{32},\ {u}_{31},\ {u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$
$I_{15}=\left\langle{u}_{21},\ {u}_{11},\ {u}_{31}^{2}\right\rangle$
$I_{16}=\left\langle{u}_{11},\ {u}_{21}^{2}\right\rangle$
$I_{17}=\left\langle{u}_{11},\ {u}_{31}^{2},\ {u}_{22}{u}_{31}-{u}_{21}{u}_{32},\ {u}_{21}{u}_{31},\ {u}_{21}^{2}\right\rangle$
$I_{18}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11},\ {u}_{31}^{2}\right\rangle$
$I_{19}=\left\langle{u}_{11},\ {u}_{32}^{2}-{u}_{22}{u}_{33},\ {u}_{31}{u}_{32}-{u}_{21}{u}_{33},\ {u}_{31}^{2},\ {u}_{22}{u}_{31}-{u}_{21}{u}_{32},\ {u}_{21}{u}_{31},\ {u}_{21}^{2}\right\rangle$
$I_{20}=\left\langle{u}_{31},\ {u}_{22},\ {u}_{21},\ {u}_{11},\ {u}_{32}^{2}\right\rangle$
$I_{21}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11},\ {u}_{32}^{2},\ {u}_{31}{u}_{32},\ {u}_{31}^{2}\right\rangle$
Figure 1: Ideals of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$ defined for the transition system $\mathcal{T}_{\mathscr{I}}$ in Example 6.10
$I_{1}$$I_{6}$$I_{8}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{2}$$I_{7}$$I_{9}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{15}}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{3}$$\boxed{I_{16}}$$I_{8}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{4}$$I_{9}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{17}}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{18}}$$I_{12}$$I_{13}$$I_{14}$$I_{5}$$\boxed{I_{19}}$$I_{11}$$\boxed{I_{20}}$$I_{13}$$I_{14}$$\boxed{I_{21}}$$\boxed{I_{20}}$$I_{13}$$I_{14}$
Figure 2: Transition forest constructed for Example 6.10. The boxed ideals $\boxed{I}$ are the ideals not of the form $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ for any $w\in\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}$.
Example 6.11.
The family of ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(4)}\right\}$ in $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{4\times 4}]$ has 43 elements. We have used Macaulay2 to construct $\mathcal{T}_{\mathscr{I}}$ and check that it is a transition system, verifying Conjecture 6.3 when $n=4$. Our calculation gives a set $\mathcal{T}_{\mathscr{I}}$ with 86 elements. The corresponding transition forest is too large to show here, but we mention that it has vertices with $\geq 2$ children.
There is a weaker form of Conjecture 6.3 that we can check by computer in a few more cases. Let $\mathcal{I}_{n}=\{w\in S_{n}:w=w^{-1}\}$ be the subset of elements in $\mathcal{I}_{\infty}^{(n)}$ whose $n\times n$ permutation matrices have full rank. Using the same algorithm as in the preceding examples, we have been able to construct a transition system containing $\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in I_{w}\}$ for all $n\leq 6$.
7 Applications to stable limits
In this final section we study the $K$-polynomials of the varieties of the ideals $I^{\mathrm{sym}}_{w}\subset\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$ defined in (6.3). Recall that these ideals are indexed by involutions $w$ in the set $\mathcal{I}_{\infty}^{(n)}$ from (6.7).
If we assume Conjecture 6.5, then the $K$-polynomials of the determinantal ideals $I^{\mathrm{sym}}_{w}$ coincide after a change of variables with the family of orthogonal Grothendieck polynomials introduced in [28]. After reviewing some relevant definitions, we will prove that if Conjecture 6.5 holds then each of these polynomials has a well-defined “stable limit” in the ring of formal power series. This will connects the main conjectures in Section 6 to an open problem posed in [28].
7.1 Equivariant K-theory
We first recall a few facts about equivariant $K$-theory from [5, §5]. Let $R$ be a finitely generated commutative $\mathbb{K}$-algebra such that $X=\operatorname{Spec}R$ is a smooth affine variety. Suppose that a linear algebraic group $G$ acts on $X$ algebraically. The action of $G$ on $X$ makes $R$ a comodule for the Hopf algebra $\mathbb{K}[G]$. An algebraic $R$-module $M$ is an $(R,G)$-module if it also has the structure of a $\mathbb{K}[G]$-comodule in such a way that the multiplication map $R\otimes M\to M$ is a morphism of $\mathbb{K}[G]$-comodules.
The equivariant $K$-theory group $K_{G}(X)$ is the Grothendieck group of the category of finitely generated $(R,G)$-modules. We usually write $[M]$ for the class of $M$ in $K_{G}(X)$, but to emphasize the group $G$ we sometimes express this as $M]_{G}$. When $Z$ is a subscheme of $X$ with ideal $I(Z)$, we define $[Z]:=[R/I(Z)]$. Here are some relevant observations:
  • We can take the tensor product over $R$ of $(R,G)$-modules, but this does not yield a well-defined product on $K_{G}(X)$ because tensor products need not preserve exact sequences. Correcting this by defining $M][N]=\sum_{i\geq 0}(-1)^{i}[\mathrm{Tor}_{i}^{R}(M,N)]$ makes $K_{G}(X)$ into a commutative ring. It is a nontrivial fact that the smoothness of $X$ guarantees finiteness of the sum.
  • Suppose $j:Z\hookrightarrow X$ is the inclusion of a smooth $G$-invariant subvariety. If $M$ is an $(R,G)$-module then each $\mathrm{Tor}_{i}^{R}(R/I(Z),M)$ is an $(R/I(Z),G)$-module, so we can view $[M][Z]$ as an element of $K_{G}(Z)$. This defines a pullback homomomorphism $j^{*}:K_{G}(X)\to K_{G}(Z)$.
[…]
  • A $(\mathbb{K},G)$-module is just a representation of $G$, so $K_{G}(\operatorname{pt})$ is the Grothendieck ring $\operatorname{Rep}(G)$ of finite-dimensional $G$-representations. The pullback of the constant map $X\to\operatorname{pt}$ gives $K_{G}(X)$ the structure of a $\operatorname{Rep}(G)$-algebra.
We note one general fact for use in the next section.
Lemma 7.1.
Suppose $V=U\oplus U^{\prime}$ is a finite-dimensional representation of $G$, with $U$ and $U^{\prime}$ subrepresentations and $\alpha:U\hookrightarrow V$ the inclusion. Then $\operatorname{Rep}(G)\cong K_{G}(V)\xrightarrow{\alpha^{*}}K_{G}(U)\cong\operatorname{Rep}(G)$ is the identity map.
Proof.
Let $\pi:V\to U$ be the natural projection map. As $M]\in\operatorname{Rep}(G)$ is identified with $\mathbb{K}[U]\otimes_{\mathbb{K}}M]\in K_{G}(U)$, we have $\pi^{*}[\mathbb{K}[U]\otimes_{\mathbb{K}}M]=[\mathbb{K}[V]\otimes_{\mathbb{K}[U]}(\mathbb{K}[U]\otimes M)]=[\mathbb{K}[V]\otimes M],$ so $\pi^{*}:K_{G}(U)\to K_{G}(V)$ is the identity map. But $\alpha^{*}\pi^{*}=(\pi\circ\alpha)^{*}=\mathrm{id}$. ∎
Choose a torus $T\subseteq G$ and suppose $M$ is a representation of $T$ with a finite-dimensional weight decomposition, so that $M=\bigoplus_{\chi\in\mathrm{Hom}(T,\mathbb{K}^{\times})}M_{\chi}$ where each weight space
$M_{\chi}=\{m\in M:tm=\chi(t)m\}$
is finite-dimensional. A weight vector for $M$ is just an element of some $M_{\chi}$, and if $v\in M_{\chi}$ then we write $\operatorname{wt}(v)=\chi$. The character of $M$ is then defined to be
$\operatorname{\mathsf{char}}(M)=\sum_{\chi\in\mathrm{Hom}(T,\mathbb{K}^{\times})}\operatorname{dim}(M_{\chi})\chi.$
We view this as an element of the completion of the group ring Z[$\mathbb{Z}[\mathrm{Hom}(T,\mathbb{K}^{\times})]$ when there are infinitely many nonzero summands.
[…]
If $X=\operatorname{Spec}R$ has a $T$-action, then $R$ itself is a $T$-module. An $(R,T)$-module $M$ is then the same thing as an $R$-module that is $\mathrm{Hom}(T,\mathbb{K}^{\times})$-graded in the sense that $R_{\chi}M_{\psi}\subseteq M_{\chi\psi}$.
Definition 7.2 (See [32, Def. 8.21]).
If $V$ is a finite-dimensional representation of $T$, then the $K$-polynomial of a finitely generated $(\mathbb{K}[V],T)$-module $M$ is $\mathcal{K}(M)=\frac{\operatorname{\mathsf{char}}(M)}{\operatorname{\mathsf{char}}(\mathbb{K}[V])}.$
The $K$-polynomial $\mathcal{K}(M)$ belongs a priori to the field of formal Laurent series in $a_{1},a_{2},\ldots,a_{n}$ over $\mathbb{K}$. It is known [32, Thm. 8.20] that this series actually gives a polynomial
$\mathcal{K}(M)\in\mathbb{K}[a_{1},a_{2},\dots,a_{n}].$ (7.1)
This polynomial exactly describes the class of $M$ in $K_{T}(V)$, as we now explain. We have
[…]
where the last isomorphism is the character map. Let $\pi:V\to\{0\}$ be the constant map. By definition, $\pi^{*}(f)=[\mathbb{K}[V]\otimes_{\mathbb{K}}W]$ for a $T$-module $W$ with character $f$. Thus $\operatorname{\mathsf{char}}(\pi^{*}(f))=\operatorname{\mathsf{char}}(\mathbb{K}[V])\operatorname{\mathsf{char}}(W)$, so we can recover $f$ from $\pi^{*}(f)$ as $\mathcal{K}(\pi^{*}(f))$.
[…]
We identify $K_{T}(V)$ with the ring of Laurent polynomials Z[$\mathbb{Z}[a_{1}^{\pm 1},a_{2}^{\pm 1},\ldots,a_{n}^{\pm 1}]$ via this map.
Example 7.3.
Let $V=\mathsf{Mat}^{\mathrm{sym}}_{2\times 2}$ and suppose $Z\subset V$ is the subvariety of noninvertible matrices (denoted $X_{1}$ in Example 3.9). Relative to the action $b:M\mapsto bMb^{\top}$, the $\mathbb{K}$-vector space $V$ is a finite-dimensional representation of the group $G=\mathsf{B}_{2}$ and therefore also of its torus $T$, which consists of all invertible $2\times 2$ diagonal matrices over $\mathbb{K}$. Let $a_{i}:T\to\mathbb{K}$ for each $i\in\{1,2\}$ denote the representation $a_{i}(t)=t_{ii}$.
In the notation of Example 3.9 we have $I(Z)=\langle u_{21}^{2}-u_{11}u_{22}\rangle$ and $Z]_{T}=[\mathbb{K}[V]/I(Z)]_{T}.$ Let $\mathcal{M}$ be the set of all monomials in the variables $\{u_{11},u_{21},u_{22}\}$. These monomials are weight vectors for $T$ with weights $\operatorname{wt}(u_{ij})=a_{i}a_{j}$ and $\operatorname{wt}(m_{1}m_{2})=\operatorname{wt}(m_{1})\operatorname{wt}(m_{2})$, and we have
$\operatorname{\mathsf{char}}(\mathbb{K}[V])=\sum_{m\in\mathcal{M}}\operatorname{wt}(m).$
If we further let $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ be the subset of monomials divisible by $u_{21}^{2}$, then $\mathcal{M}\setminus\mathcal{M}^{\prime}$ is a basis of weight vectors for $\mathbb{K}[V]/I(Z)$. Therefore, we have
$\mathcal{K}(Z):=\mathcal{K}(V/I(Z))=\tfrac{\operatorname{\mathsf{char}}(\mathbb{K}[V]/I(Z))}{\operatorname{\mathsf{char}}(\mathbb{K}[V])}=\tfrac{\sum_{m\in\mathcal{M}}\operatorname{wt}(m)-\sum_{m\in\mathcal{M}^{\prime}}\operatorname{wt}(m)}{\sum_{m\in\mathcal{M}}\operatorname{wt}(m)}=1-\tfrac{\sum_{m\in\mathcal{M}^{\prime}}\operatorname{wt}(m)}{\sum_{m\in\mathcal{M}}\operatorname{wt}(m)}.$
The map $m\mapsto u_{21}^{2}m$ is a bijection $\mathcal{M}\to\mathcal{M}^{\prime}$ with $\operatorname{wt}(u_{21}m)=(a_{1}a_{2})^{2}\operatorname{wt}(m)$, so in fact
$\mathcal{K}(Z)=1-(a_{1}a_{2})^{2}.$
If $v_{1},v_{2},\ldots,v_{m}$ is a basis of weight vectors for any $T$-representation $V$, then one always has the explicit formula $\operatorname{\mathsf{char}}(\mathbb{K}[V])=\prod_{i=1}^{m}\tfrac{1}{1-\operatorname{wt}(v_{i})}$. However, as this example suggests, this computation is not really necessary for determining $K$-polynomials.
7.2 Orthogonal Grothendieck polynomials
From now on we specialize to the case when $V=\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ and $T$ is the group of invertible $n\times n$ diagonal matrices over our algebraically closed field $\mathbb{K}$, for an arbitrary positive integer $n$. Then $V$ is a finite-dimensional representation of $T$ under the action $t:M\mapsto tMt$.
As in Section 6 and Example 7.3, we identify $\mathbb{K}[V]=\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]=\mathbb{K}[u_{ij}:1\leq j\leq i\leq n]$ and let $a_{i}:T\to\mathbb{K}$ for $i\in[n]$ denote the representation $a_{i}(t)=t_{ii}$. Then each variable $u_{ij}$ is a weight vector in the $T$-module $\mathbb{K}[V]$ with weight $a_{i}a_{j}$, and we identify
[…]
Fix an involution $w\in\mathcal{I}_{\infty}^{(n)}$. Recall the variety $\mathsf{MSV}^{\mathrm{sym}}_{w}$ from Definition 6.1 and the ideal $I^{\mathrm{sym}}_{w}$ defined in (6.3). The quotient $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]/I^{\mathrm{sym}}_{w}$ is a finitely generated $(\mathbb{K}[V],T)$-module, so we may introduce the following definitions.
Definition 7.4.
For each $w\in\mathcal{I}_{\infty}^{(n)}$ let $\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}=\mathcal{K}(\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]/I^{\mathrm{sym}}_{w})$ and $\mathfrak{G}^{\textsf{O}}_{w}=[\mathsf{MSV}^{\mathrm{sym}}_{w}]_{T}.$
From the discussion in the previous section and [28, §2.4], we know that $\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}$ and $\mathfrak{G}^{\textsf{O}}_{w}$ are both elements of $\mathbb{K}[a_{1},a_{2},\dots,a_{n}]$. The polynomial $\mathfrak{G}^{\textsf{O}}_{w}$ is essentially the orthogonal Grothendieck polynomial denoted by the same notation in [28], except that here we have set a bookkeeping parameter $\beta=-1$. The substitution $\beta=-1$ loses no information since the parameter $\beta$ can be reintroduced by a simple change of variables [28, §2.2].
Recall that Conjecture 6.5 asserts in part that $I^{\mathrm{sym}}_{w}=I(\mathsf{MSV}^{\mathrm{sym}}_{w})$, which implies that $\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}=\mathfrak{G}^{\textsf{O}}_{w}$. We adopt this conjecture as a hypothesis throughout this section, but preserve the distinct notations $\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}$ and $\mathfrak{G}^{\textsf{O}}_{w}$ to clarify our arguments.
[…]
Let $x_{1},x_{2},x_{3},\dots$ be the sequence of commuting formal variables with $x_{i}=1-a_{i}$. We may then view $\mathfrak{G}^{\textsf{O}}_{w}$ (as well as $\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}$) as an element of the polynomial ring $\mathbb{K}[x_{1},x_{2},\dots,x_{n}].$ It is an open problem [28, Prob. 5.3] to show that the stable limit
$\lim_{m\to\infty}\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}$ (7.2)
converges in the sense of formal power series to an element of $\mathbb{K}\llbracket x_{1},x_{2},\dots\rrbracket$ for all $w\in\mathcal{I}_{\infty}^{(n)}$.
We prove in this section that if Conjecture 6.5 holds then this limit does exist, and in fact gives a symmetric function, meaning a power series invariant under all permutations of the $x$-variables. This convergence is presently only known to hold unconditionally for involutions $w=w^{-1}\in S_{n}$ that are vexillary in the sense of being $2143$-avoiding [28].
Example 7.5.
[…]
We report a few computations related to the simplest non-vexillary $w\in I_{n}$. Assume $\mathbb{K}$ has characteristic zero. For $m=1,2,3,4,5$ we have used Macaulay2 to calculate that
$\mathfrak{G}^{\textsf{O}}_{1^{m-1}\times 2143}=2\cdot G\hskip-0.56905ptQ_{(2)}(x_{1},x_{2},\dots,x_{m})$ $-G\hskip-0.56905ptQ_{(2,1)}(x_{1},x_{2},\dots,x_{m})$
$-G\hskip-0.56905ptQ_{(3)}(x_{1},x_{2},\dots,x_{m})$
$+G\hskip-0.56905ptQ_{(3,1)}(x_{1},x_{2},\dots,x_{m})+x_{m+1}\cdot A_{m}+x_{m+2}\cdot B_{m}$
for certain extra terms $A_{m},B_{m}\in\mathbb{K}[x_{1},x_{2},\dots,x_{m+2}]$. Here $G\hskip-0.56905ptQ_{\lambda}(x_{1},x_{2},\dots,x_{m})$ means the polynomial obtained from $G\hskip-0.56905ptQ_{\lambda}$ by setting $x_{m+1}=x_{m+2}=\dots=0$, which is zero if $m$ is smaller than the number of nonzero parts in $\lambda$.
It would follow that $\lim_{m\to\infty}\mathfrak{G}^{\textsf{O}}_{1^{m}\times 2143}=2\cdot G\hskip-0.56905ptQ_{(2)}-G\hskip-0.56905ptQ_{(2,1)}-G\hskip-0.56905ptQ_{(3)}+G\hskip-0.56905ptQ_{(3,1)}$ if this pattern were to continue. (Notice that the sign of the coefficient of each $G\hskip-0.56905ptQ_{\lambda}$ in this expression is the same as $(-1)^{|\lambda|}$. The question of whether this sign alternation is a general phenomenon is open; see [28, Prob. 5.4].) Our methods are only powerful enough to establish convergence of the stable limit (7.2) when Conjecture 6.5 holds, however, rather than any exact formulas.
Before getting to our main results, we need to derive several technical lemmas.
Lemma 7.6.
Let $\iota:\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\to\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}$ be the inclusion $A\mapsto 1\times A$. Then $\iota^{*}(I^{\mathrm{sym}}_{1\times w})=I^{\mathrm{sym}}_{w}$ where $\iota^{*}$ denotes the induced morphism $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}]\to\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$.
Proof.
For each $i,j\in[n]$ let $\mathcal{M}_{ij}(w)$ be the set of minors of size $\operatorname{rank}w_{[i][j]}+1$ inside $\mathcal{U}^{\mathrm{sym}}_{[i][j]}$, so that the ideal $I^{\mathrm{sym}}_{w}$ is generated by the set
$\mathcal{M}(w):=\bigcup_{i,j\in[n]}\mathcal{M}_{ij}(w).$ (7.3)
The effect of applying $\iota^{*}$ is to evaluate a polynomial in $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}]$ on the matrix $1\times\mathcal{U}^{\mathrm{sym}}_{[n][n]}$. Setting $\tilde{R}=\{i-1:i\in R\setminus\{1\}\}$ for a set $R\subseteq[n]$, we therefore have
$\iota^{*}(\det\mathcal{U}^{\mathrm{sym}}_{RC})=\begin{cases}0&\text{if $1\in R\mathbin{\triangle}C$ (with $\mathbin{\triangle}$ denoting symmetric difference)}\\ \det\mathcal{U}^{\mathrm{sym}}_{\tilde{R}\tilde{C}}&\text{otherwise}.\end{cases}$
Now consider the ideal $\iota^{*}(I^{\mathrm{sym}}_{1\times w})$ generated by $\bigcup_{i,j\leq n+1}\iota^{*}\mathcal{M}_{ij}(1\times w)$. We observe that:
  • If $\min(i,j)=1$, then $\operatorname{rank}(1\times w)_{[i][j]}+1=2$, so $\mathcal{M}_{ij}(1\times w)$ is empty.
  • If $i,j>1$, then $\operatorname{rank}(1\times w)_{[i][j]}+1=\operatorname{rank}w_{[i-1][j-1]}+2$, so $\iota^{*}(\mathcal{M}_{ij}(1\times w))$ is the sum of $\mathcal{M}_{i-1,j-1}(w)$ (coming from the minors involving row and column 1) and certain minors of degree $\operatorname{rank}(w_{[i-1][j-1]})+2$ (coming from the minors not involving row or column 1). However, the latter are in the ideal generated by the former, by the general fact that the ideal of size $r$ minors in a matrix contains the minors of size $r+1$.
Thus $\bigcup_{i,j\leq n+1}\iota^{*}(\mathcal{M}_{ij}(1\times w))$ and $\mathcal{M}(w)=\bigcup_{i,j\leq n}\mathcal{M}_{ij}(w)$ generate the same ideal. ∎
Below, we distinguish the tori in the algebraic groups $\mathsf{B}_{n}$ and $\mathsf{B}_{n+1}$ by writing $T_{n}$ for the former and $T_{n+1}$ for the latter. Let $t\in T_{n}$ act on $\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}$ by the matrix $\iota(t)$. Then $\iota$ is $T_{n}$-equivariant, so there is a pullback map $\iota^{*}:K_{T_{n}}(\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)})\to K_{T_{n}}(\mathsf{Mat}^{\mathrm{sym}}_{n\times n})$. The group homomorphism $\iota|_{T_{n}}:T_{n}\to T_{n+1}$ also induces a restriction homomorphism
[…]
which simply sends the $K_{T_{n+1}}$-class of a module to its $K_{T_{n}}$-class.
Lemma 7.7.
Letting $\iota^{*}$ and $\phi$ be as above, we have
$\iota^{*}(\phi(a_{i}))=\begin{cases}1&\text{if $i=1$}\\ a_{i-1}&\text{otherwise}.\end{cases}$
Proof.
[…]
Now use Lemma 7.1 to deduce the formula for $\iota^{*}(\phi(a_{i}))$. ∎
For the next lemma, we make use of the following concepts from [20, Chapter III].
Definition 7.8 (See [20, Chapter III, §9]).
Let $C$ be a smooth algebraic curve and $X$ any variety. Let $p_{C}:X\times C\to C$ and $p_{X}:X\times C\to X$ be the projections. A collection of subschemes
$\{Z_{t}\subseteq X:t\in C\}$
is a flat family (over $C$) if there is a closed subscheme $Z^{\prime}\subseteq X\times C$ for which the restriction $p_{C}|_{Z^{\prime}}$ is a flat morphism and $Z_{t}=p_{X}(p_{C}^{-1}(t))$ for all $t\in C$.
Recall that $G$ is a linear algebraic group.
Lemma 7.9.
[…]
Proof.
[…]
$p_{C}|_{Z^{\prime}}^{*}([t])=[p_{C}|_{Z^{\prime}}^{-1}(t)]=[Z_{t}\times\{t\}]\in K_{G}(Z^{\prime}).$
[…]
Now consider the commutative diagram
[…]
where $\tilde{p}_{X}$ is the restriction of $p_{X}$ to $Z_{t}\times\{t\}$ and $\alpha$ is inclusion. By [20, Ch. III, Prop. 9.3], the higher direct image functor (as defined [20, Ch. III, §8]) of the pushforward of $p_{X}$ satisfies
$R^{i}(p_{X})_{*}(\mathcal{O}_{Z_{t}\times\{t\}})=R^{i}(\tilde{p}_{X})_{*}(\alpha^{*}\mathcal{O}_{Z_{t}\times\{t\}}).$
Since $\tilde{p}_{X}$ has 0-dimensional fibers, [20, Ch. III, Cor. 11.2] shows that the right-hand side is 0 for $i>0$. Therefore the pushforward $(p_{X})_{*}:K_{G}(X\times C)\to K_{G}(X)$ sends $\mathcal{O}_{Z_{t}\times\{t\}}]$ to $(p_{X})_{*}\mathcal{O}_{Z_{t}\times\{t\}}]$, i.e., sends the class $Z_{t}\times\{t\}]$ to $p_{X}(Z_{t}\times\{t\})]=[Z_{t}]$. Therefore $Z_{t}]$ is also independent of $t$. ∎
Now suppose $Z\subseteq X\times(C\setminus\{c\})$ is a subscheme where $C$ is a smooth curve and $c\in C$. Then there is a corresponding family of subschemes $\{Z_{t}:t\in C\setminus\{c\}\}$, where
$Z_{t}:=\pi_{X}(Z\cap\pi_{C}^{-1}(t))$
is the fiber of $Z$ over $t$. We define $\lim_{t\to c}Z_{t}$ to be the fiber of the closure $\overline{Z}\subseteq X\times C$ over $c$. We sometimes denote this limit by
$Z_{c}:=\lim_{t\to c}Z_{t}.$ (7.4)
It turns out [20, Ch. III, Prop. 9.8] that if $\{Z_{t}:t\in C\setminus\{c\}\}$ is a flat family over $C\setminus\{c\}$, then the enlarged family $\{Z_{t}:t\in C\}$ is also flat over $C$.
Define $p:\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}\to\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ by the formula $p(A)=A_{[2,n+1][2,n+1]}$. This is a left inverse of the map $\iota$ from Lemma 7.6. For $t\in\mathbb{K}^{\times}$, define $e_{t}:\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}\to\mathsf{Mat}^{\mathrm{sym}}_{(n+1)\times(n+1)}$ by
$e_{t}(A)_{ij}=\begin{cases}tA_{11}&\text{if $(i,j)=(1,1)$}\\ A_{ij}&\text{otherwise}.\end{cases}$
Lemma 7.10.
If $w\in\mathcal{I}_{\infty}^{(n)}$ then $\lim_{t\to 0}e_{t}(\mathsf{MSV}^{\mathrm{sym}}_{1\times w})$ contains $p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w})$.
Proof.
[…]
$f(t)=\begin{cases}at^{-1}&\text{if $a\neq 0$}\\ 1&\text{if $a=0$}\end{cases}\qquad\text{and}\qquad M(t)=\begin{bmatrix}f(t)&v^{\top}\\ v&A^{\prime}+f(t)^{-1}vv^{\top}\end{bmatrix}$.
We claim that $M(t)\in\mathsf{MSV}^{\mathrm{sym}}_{1\times w}$. Choose indices $i,j\in[n+1]$. We must check that
$\operatorname{rank}M(t)_{[i][j]}\leq\operatorname{rank}(1\times w)_{[i][j]}.$
This is trivial if $i$ or $j$ is $1$, so assume $i,j>1$. Then we have
$M(t)_{[i][j]}=\begin{bmatrix}f(t)&y^{\top}\\ x&A^{\prime}_{[i-1][j-1]}+f(t)^{-1}xy^{\top}\end{bmatrix}$
for some $x,y\in\mathbb{K}^{n}$. But this matrix is equivalent by row operations to $\left[\begin{smallmatrix}f(t)&y^{\top}\\ 0&A^{\prime}_{[i-1][j-1]}\end{smallmatrix}\right]$, so
$\operatorname{rank}M(t)_{[i][j]}=\operatorname{rank}\left[\begin{smallmatrix}f(t)&y^{\top}\\ 0&A^{\prime}_{[i-1][j-1]}\end{smallmatrix}\right]$ $\leq 1+\operatorname{rank}A^{\prime}_{[i-1][j-1]}$
$\leq 1+\operatorname{rank}w_{[i-1][j-1]}\qquad\text{(since $A^{\prime}\in\mathsf{MSV}^{\mathrm{sym}}_{w}$)}$
$=\operatorname{rank}\,(1\times w)_{[i][j]}.$
Thus $M(t)\in\mathsf{MSV}^{\mathrm{sym}}_{1\times w}$. Since $\lim_{t\to 0}tf(t)=a$, we have $\lim_{t\to 0}e_{t}M(t)=A$, so we have shown that $\lim_{t\to 0}e_{t}\mathsf{MSV}^{\mathrm{sym}}_{1\times w}$ contains $p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w})$ as desired. ∎
Lemma 7.11.
Suppose Conjecture 6.5 holds and $w\in\mathcal{I}_{\infty}^{(n)}$. Then $\mathfrak{G}^{\textsf{O}}_{w}$ is obtained from the polynomial $\mathfrak{G}^{\textsf{O}}_{1\times w}\in\mathbb{K}[a_{1},a_{2},\dots,a_{n+1}]$ by substituting $a_{1}\mapsto 0$ and $a_{i+1}\mapsto a_{i}$ for all $i\in[n]$.
Proof.
Recall the definition of $\mathcal{M}(w)$ from (7.3). For $t\neq 0$, let $V_{t}$ be the variety $e_{t}(\mathsf{MSV}^{\mathrm{sym}}_{1\times w})$, and set $V_{0}=\lim_{t\to 0}V_{t}$. The ideal of $V_{t}$ is $(e_{t}^{*})^{-1}I(\mathsf{MSV}^{\mathrm{sym}}_{1\times w})$, which, assuming Conjecture 6.5, equals $(e_{t}^{*})^{-1}I^{\mathrm{sym}}_{1\times w}$. The latter is generated by the minors $\mathcal{M}(1\times w)$ except that all monomials not divisible by $u_{11}$ get multiplied by $t$.
Setting $t=0$, we see that $I(V_{0})$ contains $u_{11}p^{*}\mathcal{M}(w)$. Hence $V_{0}$ is contained in the union
$\{u_{11}=0\}\cup p^{-1}(V(I^{\mathrm{sym}}_{w}))=\{u_{11}=0\}\cup p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w}),$
again using Conjecture 6.5. On the other hand, $V_{0}$ contains $p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w})$ by Lemma 7.10. Therefore we can write $V_{0}=Z\cup p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w})$ where $Z$ is a closed subscheme of $\{u_{11}=0\}$.
Now consider the $K$-class
$V_{0}]=[Z]+p^{*}[\mathsf{MSV}^{\mathrm{sym}}_{w}]-[Z\cap p^{-1}(\mathsf{MSV}^{\mathrm{sym}}_{w})].$
Any closed subscheme of $\{u_{11}=0\}$ has zero $K_{T_{n}}$-class since $T_{n}$ acts trivially on $u_{11}$. Therefore
$V_{0}]=p^{*}[\mathsf{MSV}^{\mathrm{sym}}_{w}].$
The subschemes $e_{t}V(I^{\mathrm{sym}}_{1\times w})$ are all isomorphic for $t\neq 0$, so they form a flat family over $\mathbb{A}^{1}\setminus\{0\}$ by [20, Ch. III, Thm. 9.9]. Lemma 7.9 therefore implies that
$\mathsf{MSV}^{\mathrm{sym}}_{1\times w}]=[V_{0}]=p^{*}[\mathsf{MSV}^{\mathrm{sym}}_{w}].$
As $p\circ\iota=\mathrm{id}$, we conclude that $\iota^{*}[\mathsf{MSV}^{\mathrm{sym}}_{1\times w}]=[\mathsf{MSV}^{\mathrm{sym}}_{w}]$. By Lemma 7.7 this completes the proof. ∎
Lemma 7.12.
If $w\in\mathcal{I}_{\infty}^{(n)}$ then the polynomial $\widetilde{\mathfrak{G}}^{\textsf{O}}_{1^{m}\times w}\in\mathbb{K}[a_{1},a_{2},\dots,a_{m+n}]$ is symmetric in the variables $a_{1},a_{2},\ldots,a_{m}$.
Proof.
Identify $\sigma\in S_{m}$ with the permutation matrix $\sigma\times 1^{n}$, so that $S_{m}$ acts on $\mathsf{Mat}^{\mathrm{sym}}_{m+n\times m+n}$ by conjugation. The induced map on the coordinate ring $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{m+n\times m+n}]$ sends $u_{ij}$ to $u_{\sigma(i),\sigma(j)}$. For each $i,j$, let $\mathcal{M}_{ij}$ be the set of minors of size $\operatorname{rank}(1^{m}\times w)_{[i][j]}+1$ inside $\mathcal{U}^{\mathrm{sym}}_{[i][j]}$. Recall that $I^{\mathrm{sym}}_{1^{m}\times w}$ is generated by $\bigcup_{i,j}\mathcal{M}_{ij}$. Now, we observe the following properties:
  • If $\min(i,j)\leq m$, then $\operatorname{rank}({1^{m}\times w})_{[i][j]}+1=\min(i,j)+1$, so $\mathcal{M}_{ij}$ is empty.
  • If $\sigma$ is a permutation fixing every $k>\max(i,j)$, then $\sigma$ maps $\mathcal{M}_{ij}$ to itself; in particular, this holds if $\sigma\in S_{m}$ and $i,j>m$.
It follows that $\sigma$ maps $\bigcup_{i,j}\mathcal{M}_{ij}$ to itself, and hence also maps $I^{\mathrm{sym}}_{1^{m}\times w}$ to itself. The action of $\sigma$ on $\mathsf{Mat}^{\mathrm{sym}}_{m+n\times m+n}$ induces an action on $K_{T_{m+n}}(\mathsf{Mat}^{\mathrm{sym}}_{m+n\times m+n})$ permuting the variables $a_{1},\dots,a_{m}$, and $\sigma I^{\mathrm{sym}}_{1^{m}\times w}=I^{\mathrm{sym}}_{1^{m}\times w}$ implies that $\widetilde{\mathfrak{G}}^{\textsf{O}}_{1^{m}\times w}=[V(I^{\mathrm{sym}}_{1^{m}\times w})]\in\mathbb{K}[a_{1},a_{2},\dots,a_{m+n}]$ is also fixed by $\sigma$. ∎
Recall that $x_{i}=1-a_{i}$ and that we consider $\mathfrak{G}^{\textsf{O}}_{w}$ as a polynomial in $x_{i}$.
Theorem 7.13.
Assume that Conjecture 6.5 holds. Choose any element $w\in\mathcal{I}_{\infty}^{(n)}$. Then the stable limit $\lim_{m\to\infty}\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}$ converges to a symmetric element of $\mathbb{K}\llbracket x_{1},x_{2},x_{3},\dots,\rrbracket$.
Proof.
Write $\mathfrak{G}^{\textsf{O}}_{1\times w}=\mathfrak{G}^{\textsf{O}}_{1\times w}(x_{1},x_{2},\ldots,x_{n+1})$ and $\mathfrak{G}^{\textsf{O}}_{w}=\mathfrak{G}^{\textsf{O}}_{w}(x_{1},x_{2},\ldots,x_{n})$ as polynomials in the $x$-variables. Then Lemma 7.11 shows that $\mathfrak{G}^{\textsf{O}}_{1\times w}(0,x_{1},\ldots,x_{n})=\mathfrak{G}^{\textsf{O}}_{w}(x_{1},\ldots,x_{n})$. Now if $x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}$ is some monomial with $d<m$ then
$[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}(x_{1},\ldots,x_{m+n})$ $=[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}(x_{1},\ldots,x_{m-1},0,\ldots,0)$
$=[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}(0,x_{1},\ldots,x_{m-1},0,0,\ldots,0)$
$=[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{m-1}\times w}(x_{1},\ldots,x_{m-1},0,0,\ldots,0),$
using Lemma 7.12 for the second equality (plus the assumption that $\mathfrak{G}^{\textsf{O}}_{w}=\widetilde{\mathfrak{G}}^{\textsf{O}}_{w}$). By induction
$\lim_{m\to\infty}[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}=[x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}}]\mathfrak{G}^{\textsf{O}}_{1^{d}\times w}(x_{1},\ldots,x_{d},0,\ldots,0).$
Thus the sequence of coefficients of any fixed monomial in $\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}$ is eventually constant as $m\to\infty$, so $\lim_{m\to\infty}\mathfrak{G}^{\textsf{O}}_{1^{m}\times w}$ converges to a formal power series, which is symmetric by Lemma 7.12. ∎